### v1

# Unreal Robotics Lab: A High-Fidelity Robotics Simulator with Advanced Physics and Rendering

Abstract
High-fidelity simulation is essential for robotics research, enabling safe and efficient testing of perception, control, and navigation algorithms. However, achieving both photorealistic rendering and accurate physics modeling remains a challenge. This paper presents a novel simulation framework–the Unreal Robotics Lab (URL) that integrates the Unreal Engine’s advanced rendering capabilities with MuJoCo’s high-precision physics simulation. Our approach enables realistic robotic perception while maintaining accurate physical interactions, facilitating benchmarking and dataset generation for vision-based robotics applications. The system supports complex environmental effects, such as smoke, fire, and water dynamics, which are critical for evaluating robotic performance under adverse conditions. We benchmark visual navigation and SLAM methods within our framework, demonstrating its utility for testing real-world robustness in controlled yet diverse scenarios. By bridging the gap between physics accuracy and photorealistic rendering, our framework provides a powerful tool for advancing robotics research and sim-to-real transfer.
I INTRODUCTION
Simulation plays a crucial role in robotics research, enabling safe, efficient, and cost-effective development, testing, and validation of robotic systems before real-world deployment. High-fidelity simulation environments provide realistic perception and physics interactions, bridging the gap between simulation and reality. However, achieving both photorealism and accurate physics simultaneously remains a challenge due to computational constraints and the complexity of real-world dynamics [1].
This paper presents a novel simulation pipeline that combines the rendering capabilities of Unreal Engine [2] with the advanced physics simulation of MuJoCo [3] to create photorealistic environments for robotics applications. By integrating these two frameworks, our approach delivers high-quality simulation results suitable for tasks such as perception training, dataset generation, and benchmarking.
A key advantage of this pipeline is its ability to generate complex and dynamic environmental effects—such as smoke, fire, and water waves—which are crucial for training robots to handle diverse and unpredictable conditions. These effects enable the simulation of rare events that are difficult to capture in real-world datasets or reproduce in existing simulators. As a result, they can be used to benchmark vision-based policies under out-of-distribution conditions and generate training data for visual-based navigation.
We present a novel, open-source simulation framework - the Unreal Robotics Lab (URL) that integrates Unreal Engine for photorealistic rendering with MuJoCo for accurate physics simulation. Unlike existing simulators focused on specific domains (e.g., autonomous driving or aerial robotics) or embodied-ai for limited robots, our system is designed for general-purpose robotic applications, supporting quadrupeds, manipulators, UAVs, and more. Additionally, we incorporate CoACD-based convex decomposition [4] for improved non-convex object representation. Our key contributions are:
  • High-Fidelity Physics with Photorealistic Rendering: Synchronization of MuJoCo’s physics with Unreal Engine’s advanced rendering (Lumen, Nanite, Niagara) enables realistic perception-based training and benchmarking.
  • Generalized Benchmarking for Diverse Conditions: A structured benchmarking system evaluates robotic perception and control under rare conditions (e.g., smoke, fog, dynamic lighting, dynamic movement).
  • Scene Generation: A GPT-based module automates diverse environment creation, reducing manual setup and enabling adaptive scene variations.
  • Comprehensive Sensor and Control Framework: Built-in support for cameras, IMUs, force sensors, and controllers for various robot types, with full Robot Operating System (ROS) compatibility.
  • Open-Source and Extensible: The system will be released as open-source, supporting modular extensions for new controllers, planners, and robotic platforms.
The subsequent sections are structured as follows. Sec. II provides an overview of the relevant literature, discussing previous works in the field, while Sec. III outlines the proposed method in detail. Sec. IV presents the experimental setup, including both simulation and real robot experiments. Finally, in Sec. V, we present the conclusions and discuss future
II RELATED WORK
Robotic simulation has advanced significantly in recent years, with various frameworks addressing key challenges such as physics accuracy, photorealistic rendering, and real-time performance. This section reviews related work in three main areas: robotics-focused simulation platforms, comparisons of physics engines, and benchmarking SLAM and navigation under adverse visual conditions.
Several other game engines and simulation platforms offer photorealistic rendering [5, 6, 7], but Unreal Engine provides distinct advantages through Lumen global illumination and Nanite virtualized geometry. Unlike other simulators that rely on hardware-intensive ray tracing, Lumen [8] enables real-time global illumination and reflections without requiring precomputed lighting, making it well-suited for dynamic environments where lighting conditions change.
Furthermore, Nanite’s automatic level of detail (LOD) [9] system allows highly detailed assets with millions of triangles to be rendered efficiently, without affecting performance. This is particularly beneficial for large-scale or high-fidelity indoor environments where maintaining real-time performance is critical. Beyond rendering, Unreal Engine offers a flexible and extensible ecosystem, supporting third-party plugins, visual scripting, AI-driven NPCs, and cross-platform deployment, making it highly adaptable for a variety of robotic simulation tasks. These features ensure that simulations can achieve both high physical accuracy and photorealistic perception, making them an ideal choice for vision-based robotics research.
II-A Robotics-Focused Simulation Platforms
Robotic simulation has made considerable progress, with numerous frameworks addressing challenges in physics accuracy, rendering quality, and real-time performance. This section reviews related work in three primary areas: robotics simulation platforms, comparisons of physics engines, and benchmarking of SLAM and navigation under adverse visual conditions.
Several simulation platforms have been developed for robotic perception, control, and reinforcement learning, each with different trade-offs in realism and efficiency. Gazebo [10] is a widely-used open-source simulator that supports multiple physics engines; ODE [10], Simbody [11], (Py)Bullet [12], and DART [13], although its rendering capabilities lag behind modern game-engine-based simulators. MuJoCo [3] excels in efficient and accurate physics modeling for robotic control but lacks built-in photorealistic rendering. GPU-accelerated platforms such as Isaac Orbit [5] and Isaac Gym [14] enable large-scale reinforcement learning, prioritizing speed over advanced rendering. Intel SPEAR [15] provides photorealistic indoor environments suitable for embodied AI research. CARLA [7] and AirSim [16] leverage Unreal Engine for photorealistic urban and aerial simulations, with a primary focus on autonomous driving and drone applications. CARLA, in particular, emphasizes high-level decision-making, such as traffic light interactions and multi-agent traffic flow, rather than fine-grained vehicle dynamics like tire skidding. However, both simulators remain largely domain-specific, limiting their utility for broader robotics applications beyond driving and aerial scenarios. RoboSuite [17] and RoboCasa [18] emphasize robotic manipulation using MuJoCo for precise physics but offer limited environmental complexity. MetaUrban [19] simulates realistic urban environments with dynamic traffic, beneficial for multi-agent interactions.
Although most simulators prioritize either physics accuracy or visual realism, few achieve both. Our approach bridges this gap by integrating MuJoCo for precise physics with Unreal Engine for state-of-the-art rendering, creating a high-fidelity simulation pipeline for diverse robotic applications.
II-B Physics Engines for Robotics Simulation
Physics engines are essential for realistic robotic simulation, particularly for modeling contact, soft-body dynamics, and fluid interactions. Various engines have been evaluated in terms of accuracy, efficiency, and scalability, with notable differences in their suitability for robotics [20]. MuJoCo [3] excels in soft-contact modeling, accuracy, and stability, making it ideal for complex robotic tasks such as grasping and locomotion. However, it lacks built-in high-fidelity rendering. ODE (Open Dynamics Engine) [10], commonly used in Gazebo, faces stability issues in high-DOF tasks and requires extensive parameter tuning. Bullet [12] is popular for reinforcement learning and manipulation, but its precision lags behind MuJoCo.
Unreal Engine’s Chaos and Havok [21] are optimized for game physics and real-time effects, excelling in destruction, fluid, and particle-based simulations. However, they are not designed for robotic applications that need soft-contact modeling or high-precision joint dynamics. NVIDIA PhysX [22] offers GPU-accelerated simulation suitable for fast reinforcement learning, but its omission of Coriolis forces makes it unsuitable for high-accuracy control.
Our approach combines MuJoCo’s precise physics with Unreal Engine’s advanced rendering, achieving both photorealism and accurate simulation for robotic applications.
II-C Benchmarking SLAM and Navigation Under Adverse Visual Conditions
Evaluating SLAM and navigation algorithms in diverse environments is challenging, as many existing datasets capture structured scenarios under ideal conditions, making it difficult to test performance under adverse effects such as smoke, fog or dynamic lighting.
Envodat [23] provides a dataset tailored for challenging conditions, but real-world data collection remains limited in scale and diversity. Our simulation pipeline fills this gap by synthetically generating rare-event scenarios, enabling benchmarking and training of vision-based navigation policies in controlled, repeatable settings.
Current simulation platforms often emphasize either high-precision physics (e.g., MuJoCo, Isaac Gym) or photorealistic rendering (e.g., Unreal-based simulators such as Carla and AirSim). A recent simulator, SPEAR [15], integrates MuJoCo with Unreal Engine but primarily focuses on providing a Python-based OpenAI Gym environment, pre-built scenes, limited robot support, and task-specific benchmarks. In contrast, our approach integrates MuJoCo’s precise physics with Unreal Engine’s advanced rendering in a flexible and generalized simulation environment. Our framework allows users to freely configure robots and tasks, and utilizes ROS to treat simulated interactions as closely as possible to real-world scenarios. Furthermore, by leveraging Unreal’s Niagara Particle and Fluid effects, we enable simulation of adverse environmental conditions such as smoke, fire, and water waves, extending traditional benchmarking capabilities and providing robust tools for evaluating robotic perception and navigation under challenging scenarios.
III Simulator setup
III-A Overview
Our simulator (URL) integrates MuJoCo and Unreal Engine to provide a comprehensive simulation environment that closely mimics real-world interactions. Robots can be effortlessly integrated by adding corresponding MuJoCo XML descriptions and associated Unreal Engine assets. The system supports standard robot configurations and sensors, including cameras, IMUs, and force sensors, ensuring wide compatibility. The overall system can be seen in Figure 2. The system is comprised of both External components and the internal simulator. The system has a SimManager that is modular and manages submodules comprised of a Benchmark, Replay, and SceneGen modules. It manages all references to the UE data and is developed as a plugin for UE that can be added to any project.
For communication, our simulator leverages ROS to enable seamless interaction with existing robotics frameworks. Furthermore, faster custom communication methods are being explored through ZeroMQ [24] for improved real-time performance. We treat the simulation environment as a surrogate real-world, facilitating the testing and validation of robotic behaviors and perception systems in realistic scenarios.
III-B Physics Manager
Our system integrates MuJoCo and CoACD [4] within the Unreal Engine (UE) for accurate physics simulation and mesh decomposition. These systems are currently compiled into the UE project, and extend a custom interface. To represent objects in the scene, we introduce an intermediate representation, termed PhysicsObject, which encapsulates both geometric and state information. This allows us to efficiently manage and transform scene elements between UE and MuJoCo while ensuring that transformations and dynamic interactions are correctly synchronized.
Objects within the scene are categorized into predefined types. Specifically, assets are tagged under one of the following categories: EP_STATIC, EP_STATIC_COMPLEX, EP_DYNAMIC, EP_DYNAMIC_COMPLEX, and EP_PRIMITIVE_(TYPE). Static objects remain unchanged throughout the simulation, whereas dynamic objects are actively tracked, with their transformations updated asynchronously within UE. The distinction between simple and complex dynamic objects determines whether the object undergoes convex decomposition by CoACD. CoACD decomposes meshes into a series of convex hulls, whereas the regular mujoco compile will generate only one convex hull per object. This allows CoACD to represent non-convex geometry and is recommended by MuJoCo [20]. This classification is recorded within the corresponding PhysicsObject to guide further processing.
Figure 1: Unreal Engine Landscape conversion to MuJoCo HeightField.
Once asset classification is complete, mesh-based assets are exported as OBJ files, ensuring compatibility with MuJoCo. Additionally, if landscapes are present in the scene, they are scanned and converted into height-fields to be used within the simulation. This process uses a custom collision channel on the landscape, then, utilizing an autogenerated bounding box, we perform multithreaded raycasts to determine the heights to use. This bounding-box limits the raycasts to within this area only at a predefined resolution. This prevents the whole landscape from being scanned. Figure 1 demonstrates the output of this process. This method was chosen because landscapes can be very large, and the user may not be interested in converting the entire landscape for physics. It also has the added benefit of being able to model small clutter and decorations in the heightfield itself. This process enables the seamless reconstruction of large-scale terrains within MuJoCo, preserving critical environmental features necessary for accurate physics interactions.
Figure 2: Unreal - Mujoco Simulator system diagram
Following asset processing, an XML scene description is automatically generated to initialize MuJoCo. This XML encapsulates all relevant object data, including mesh file references, physics properties, and hierarchical relationships. Special attention is paid to robotic assets that are detected within the MuJoCo environment and mapped back to their corresponding UE entities. During this process, cameras and sensors specified in the XML are instantiated within UE to ensure consistency across both environments. Sensor elements are represented as sites within MuJoCo, allowing for precise placement and interaction within the physics simulation.
The physics simulation operates asynchronously, running within a separate execution thread to maximize performance. Transformation updates are also managed in a dedicated asynchronous thread within UE to maintain real-time synchronization with MuJoCo. External forces applied within MuJoCo can be propagated to UE entities, allowing for bidirectional interaction between the simulation and the rendered environment. UE physics can still be used for objects not managed by MuJoCo.
To improve computational efficiency, assets that remain unchanged between simulation runs are cached, preventing unnecessary recomputation and redundant file exports. This caching mechanism ensures that static elements do not incur additional processing overhead, allowing the system to focus on dynamic entities that require continuous state updates. The final XML scene description, along with the automatically generated mesh assets, enables direct execution within MuJoCo as a standalone simulation. A supporting script is provided to facilitate this process, allowing the same environment to be used independently of UE when required.
III-C Benchmark System
The benchmarking system automates the recording and management of evaluation metrics throughout the simulation. At its core, it consists of a MetricManager and a BaseMetric class, providing a structured and extensible framework to track performance indicators.
The BaseMetric class defines a fundamental interface for all metrics, encapsulating a two-dimensional float array that serves as a storage mechanism for time-series data. This class provides essential methods, including initialization, resetting, and exporting data to CSV format. These methods ensure that each metric implementation adheres to a consistent lifecycle, simplifying the process of integrating new metrics into the system. Derived metrics can override these base methods to implement domain-specific logic while still benefiting from the structured data storage and export mechanisms.
The MetricManager is responsible for the automatic discovery and registration of all metric implementations. It dynamically populates a list of available metrics by scanning for corresponding class definitions, allowing modular and scalable metric tracking. Users can enable or disable specific metrics through a dedicated user interface, ensuring flexibility in selecting relevant performance indicators for a given experiment.
By default, the system tracks core metrics, including: DistanceToGoal, which tracks the remaining distance from an entity to a predefined target location; Collisions, which records contact events with obstacles or other entities; TimeToGoal, which measures the time required to reach the destination; and GlobalPose, which logs the absolute position and orientation of an entity over time.
All metrics are recorded at discrete intervals and automatically saved as CSV files within designated experiment folders, ensuring organized and accessible results. The system also features grace periods (delaying data recording for $n$ seconds) and timeouts (terminating recordings after $n$ seconds), improving efficiency and reproducibility across simulation runs.
III-D Scene Generation
Figure 3: Example of room generated using the experimental Scene Generation system leveraging GPT API.
Our simulator features an experimental scene generation module designed primarily for indoor environments, leveraging OpenAI’s ChatGPT [25] API. When an instance of the ASceneGenerator Actor is present and assigned a RoomType, the system automatically generates object placements based on the specified scene type (Fig. 3). The module queries ChatGPT with a structured prompt containing a list of available assets and receives a JSON-formatted response specifying object positions. These placements are then parsed, and the corresponding assets are instantiated in the environment.
To ensure spatial coherence, the system detects overlapping objects using Unreal Engine’s physics system. If overlaps occur, an iterative refinement loop provides feedback to ChatGPT for updated placement recommendations, continuing until all collisions are resolved or a predefined iteration limit is reached. This method enables efficient, collision-free scene generation, producing diverse and structured environments suitable for robotic experimentation.
III-E Replay
Our system features an automated recording and replay mechanism for sensor and body pose data, enabling efficient data collection and experimentation. It uses custom structs to track only essential information –pose data, sensor readings, and timestamps– storing themin a queue for minimal overhead and precise sequence reproduction. Data streams are managed through separate JSON files, all of which can be dynamically saved and loaded at runtime. By publishing replayed data on the same ROS topics as live data, the system integrates seamlessly with existing pipelines. This allows researchers to collect data once and replay it in visually modified environments, supporting domain adaptation, robustness testing, and dataset augmentation for machine learning and benchmarking.
III-F Extended Features
Our system leverages Unreal Engine’s advanced rendering and simulation capabilities, surpassing traditional robotics simulators like MuJoCo [3], Gazebo [10], and Isaac Sim [5]. Key features include real-time global illumination (Lumen), high-detail mesh rendering (Nanite), and physically based reflections, which enhance perception-based robotics research. Procedural animations, NavMesh navigation, and built-in environmental tools enable efficient creation of dynamic simulations, while the Niagara particle system facilitates testing under adverse visual conditions (e.g., smoke, fog). Additionally, packaged standalone deployments improve reproducibility and scalability, addressing limitations of conventional simulators.
III-G Example Deployment of Robots, Controllers, and Visual Planners
The system provides a unified framework where robots, controllers, and planners operate as in real-world deployments. All components communicate through ROS, ensuring seamless integration with external systems.
III-G1 Example Robot Integrations
The framework supports quadrupeds, quadruped mobile manipulators, UAVs, and robotic manipulators, enabling diverse experimental setups for locomotion, aerial navigation, and manipulation.
III-G2 Controllers
The system includes joint controllers based on position and torque, as well as more advanced controllers such as Walk-These-Ways [26] and VBC [27]. UAVs use a PID-based controller, whereas manipulators rely on inverse kinematics (IK) methods such as Mink [28]. Since controllers operate within the ROS ecosystem, new control strategies can be easily integrated.
III-G3 Planners
Motion planning is handled using learning-based and optimization-based planners such as ViNT [29], GNM [30], and NoMAD [31]. The ROS-based architecture of the framework ensures compatibility with additional external planners.
III-G4 Visual SLAM
The system includes visual SLAM techniques such as OrbSLAM2 [32], OrbSLAM3 [33], and MASt3R-SLAM [34], enabling real-time localization and mapping. These methods integrate directly into the framework, enabling realistic perception and mapping in simulated environments.
IV Experiments
IV-A Simulation and Real-World Comparision
To evaluate the alignment between simulated and real-world visual inputs for image encoders, we employ Grad-CAM, EigenCAM, cosine similarity and KL Divergence. These methods assess both spatial attention and feature distribution similarity.
Grad-CAM computes the gradient of the class score $y^{c}$ with respect to feature map activations:
$\alpha_{k}^{c}=\frac{1}{Z}\sum_{i}\sum_{j}\frac{\partial y^{c}}{\partial A^{k}_{ij}},\quad L^{c}=\text{ReLU}\left(\sum_{k}\alpha_{k}^{c}A^{k}\right)$ (1)
EigenCAM generates class-agnostic saliency maps via principal component analysis (PCA):
$C=\frac{1}{Z}\sum_{i,j}(A^{k}_{ij}-\mu)(A^{k}_{ij}-\mu)^{T},\quad L_{\text{EigenCAM}}=v_{1}A^{k}$ (2)
where $\mu$ is the mean activation and $v_{1}$ is the first principal component.
Figure 4: Comparison of real-world (left) and simulated (right) images, along with their corresponding Grad-CAM (CLIP) and EigenCAM (EfficientNet-B0) heatmaps. The highlighted class label is ”cone”.
Cosine Similarity measures feature alignment between real and simulated embeddings:
$S_{\text{cosine}}(P,Q)=\frac{P\cdot Q}{\|P\|\|Q\|}$ (3)
where $P$ and $Q$ are feature vectors. Higher values indicate stronger similarity.
KL Divergence quantifies distribution differences:
$D_{\text{KL}}(P\|Q)=\sum_{i}P(i)\log\frac{P(i)}{Q(i)}$ (4)
Lower values indicate closer alignment.
To compare models trained on real-world data, we use CLIP [35] and the EfficientNet-B0 [36] visual encoder used in ViNT [29]. We first capture an image from the real world and recreate it within our simulator. Then both images are processed using Grad-CAM amd Eigen-Cam to generate attention heatmaps, as shown in Figure 4.
In both cases, the models primarily attend to the traffic cone in the center of the hallway as well as the end of the corridor, suggesting that the most salient features are preserved between the real and simulated environments. This alignment indicates that the models are focusing on the same high-level structures regardless of domain, reinforcing the realism of the simulated scene in terms of feature representation. The similarity in attention patterns also suggests that key objects remain distinguishable across domains, which is critical for transfer learning and sim-to-real applications.
To further quantify similarity, we compute KL divergence (0.4679) and cosine similarity (0.7157) between real and simulated feature distributions. Higher cosine similarity and lower KL divergence suggest closer alignment. While we did not aim to replicate real-world scenes with identical assets, opting instead for a simulation that remains visually representative of real conditions. Although the assets in our simulation differ from those in the real environment, our findings, together with GradCAM and EigenCAM comparisons, indicate that the simulation is visually representative enough for our evaluation purposes.
IV-B Visual Navigation and SLAM Benchmarking
Figure 5: Examples of the different adversity levels for the House Environment.
we benchmark standard methods to demonstrate simulator realism. We evaluated visual navigation (Experiment 1) with GNM [30] and ViNT [29], and visual SLAM (Experiment 2) with OrbSLAM2 [32], OrbSLAM3 [33], and MASt3R-SLAM [34]. Each method was run 5 times per condition, averaging results across trials. Both experiments used the Unitree Go1 quadruped with its onboard camera and the Walk-These-Ways low-level controller for realistic deployment.
The benchmarks were conducted in two distinct environments—warehouse and residential house—under three adversity levels: none, minor and major, classified by human judgment. UE’s Niagara system simulated environmental effects (smoke, rain, pollution, snow), while dynamic lighting and NPC movements were added for major adversity. Figure LABEL:fig:example_photo_realistic_render shows example environments and Figure 5 shows examples of adversity levels.
IV-B1 Experiment 1: Visual Navigation
For visual navigation benchmarking, the robot was initialized at identical starting locations across environments and adversity levels. Each method had a fixed duration to reach the goal using a pre-built image-based topological map. The warehouse and residential house were used for this experiment. Methods were not fine-tuned, relying on a simplified topological map approach from the authors’ original repository [31], which may limit performance [37].
Success Weighted by Collision (SC). Following [38], we adopt the SC metric:
$SC=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{1}{1+c_{i}},$ (5)
where $S_{i}$ is the success indicator for episode $i$, $c_{i}$ is the number of collisions, and $N$ is the total episodes. Higher SC values indicate fewer collisions in successful runs, while lower values reflect frequent or severe collisions. Alongside SC, we measure success rate, total collisions, time to goal, and goal distance.
Env Adversity Method
Success
$\uparrow$
[…]
Time
[…]
Goal
[…]
Weighted
[…]
House Major GNM 0.2 13 0 7.01 0.0154
ViNT 0 9.2 0 8.41 0
Minor GNM 1 1.2 71.59 1.99 0.5333
ViNT 0.4 2.8 0 4.75 0.09
None GNM 1 1.8 147.86 1.99 0.3667
ViNT 1 0.8 141.35 2.00 0.7333
Ware- house Major GNM 0 12.8 0 8.33 0
ViNT 0 6.4 0 7.96 0
Minor GNM 0.4 4 107.62 4.69 0.1067
ViNT 0.2 6.6 54.91 6.61 0.0250
None GNM 0.8 3.4 102.86 2.53 0.1889
ViNT 0.8 4.4 90.14 2.59 0.3667
TABLE I: Performance of different Visual navigation methods under various environments and varying adverse effects.
Table I presents the results, including Weighted SC to factor in collision penalties. State-of-the-art methods perform well under normal conditions, with ViNT generally outperforming GNM. Under minor adversity, GNM maintains a higher success rate but takes longer, while ViNT completes faster when successful. However, performance declines as environmental adversity increases (e.g., obstacles, smoke, rain), with both methods experiencing more collisions, longer times to goal, and lower SC. In severe adversity, neither method succeeds in the Warehouse, leading to zero SC. This suggests that these models struggle in extreme conditions due to limited training data, highlighting the need for training in diverse high-adversity conditions, which can be generated more effectively in simulation. The ability to create controlled but realistic adverse scenarios in a simulated environment provides a path to improve robustness in real-world deployments.
IV-B2 Experiment 2: Visual SLAM
This experiment used our replay system, where a teleoperated robot followed predefined trajectories optimized for SLAM. Initial runs recorded robot poses and sensor data, which were later replayed with identical conditions but varying visual effects. Each SLAM method was then evaluated, and performance metrics were computed.
Absolute Trajectory Error (ATE) measures the global accuracy of an estimated trajectory by computing the RMSE between estimated positions $\hat{\mathbf{p}}_{i}$ and ground truth positions $\mathbf{p}_{i}$:
$\text{ATE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(\mathbf{p}_{i}-\hat{\mathbf{p}}_{i}\right)^{2}}$ (6)
Coverage measures the ratio of the estimated trajectory length to the ground truth:
$\text{Coverage}=\frac{\sum_{i=1}^{n-1}\left\|\mathbf{p}_{i+1}-\mathbf{p}_{i}\right\|}{\sum_{i=1}^{n-1}\left\|\hat{\mathbf{p}}_{i+1}-\hat{\mathbf{p}}_{i}\right\|}$ (7)
note that since monocular SLAM estimates trajectories with arbitrary scale, Umeyama’s method [39] is applied for scale and pose correction before computing coverage. Poor tracking or alignment can result in coverage exceeding 100%.
Scaled ATE adjusts ATE for localization failures:
Scaled $\text{Scaled ATE}=\frac{\text{ATE}}{\text{Coverage}}$ (8)
Figure 6: comparision of different Visual SLAM algorithms under different visual effects
Env Adver. OrbSLAM2 Mono OrbSLAM3 Mono MASt3R-SLAM
[…]
Scaled
[…]
Cover-
$\uparrow$
[…]
Scaled
[…]
Cover-
$\uparrow$
[…]
Scaled
[…]
Cover-
$\uparrow$
Ware- house None 4.23 6.61 0.639 5.29 7.72 0.685 0.324 0.322 1.01
Minor 5.33 10.1 0.527 5.21 8.80 0.592 1.55 1.77 0.872
Major 2.96 20.2 0.147 4.29 5.87 0.731 0.252 0.417 0.605
House None 3.49 4.64 0.752 1.69 1.17 1.45 0.274 0.326 0.840
Minor 0.513 1.23 0.415 0.547 1.81 0.302 0.257 0.300 0.855
Major 0.475 2.53 0.188 2.70 11.6 0.233 0.275 0.284 0.966
TABLE II: Performance of Different VSLAM Methods under Varying Visual Effects.
From Table II and Fig. 6, Visual SLAM performance degrades with increasing visual adversity, highlighting sensitivity to challenging conditions. OrbSLAM2 struggles the most, with sharp declines in coverage and increased ATE, particularly in the warehouse environment. OrbSLAM3 performs well in clean conditions but deteriorates significantly under adversity, suggesting reduced robustness to occlusions, lighting changes, and dynamic elements. MASt3R-SLAM maintains the lowest ATE and highest coverage across conditions, demonstrating superior resilience to visual disturbances. Under severe visual adversity, all methods experience increased ATE and reduced coverage, highlighting the inherent challenges of maintaining reliable SLAM performance in dynamic or visually degraded environments. These replay-based evaluations provide a controlled, repeatable framework for assessing SLAM robustness across different levels of adversity, reinforcing the need for training and evaluation in diverse, high-adversity conditions to enhance real-world deployment reliability.
V CONCLUSION
This paper presents a high-fidelity robotics simulation framework that integrates advanced photorealistic rendering with precise physics modeling to support vision-based robotics research and benchmarking. By leveraging Unreal Engine’s rendering capabilities and MuJoCo’s accurate physics, our system enables evaluation of robotic navigation and SLAM methods under diverse and adverse environmental conditions. Experimental results confirm that our simulator provides a robust testbed for evaluating real-world robustness in controlled, repeatable scenarios.
Despite its strengths, the system has certain limitations, including the absence of direct reinforcement learning (RL) support, lack of deformable terrain simulation, limited scene generation capabilities, and a focus on general robotics rather than autonomous vehicles or high-precision drone simulations.
Future Work. To address these limitations and further enhance our simulator, we plan to further enhance the MuJoCo integration, allowing the simulator to communicate with a standalone MuJoCo process and run RL environments (MuJoCo Playground [40]), generate synthetic datasets to enhance vision-based model training, expand automated scene generation, and enable imitation learning via VR-based teleoperation.
Overall, our platform serves as a robust foundation for benchmarking and data generation, bridging the gap between simulation and real-world robotics deployment.

### v2

@@
 However, achieving both photorealistic rendering and accurate physics modeling remains a challenge.
-This paper presents a novel simulation framework–the Unreal Robotics Lab (URL) that integrates the Unreal Engine’s advanced rendering capabilities with MuJoCo’s high-precision physics simulation.
+This paper presents a novel simulation framework, the Unreal Robotics Lab (URL), that integrates the advanced rendering capabilities of the Unreal Engine with MuJoCo’s high-precision physics simulation.
 Our approach enables realistic robotic perception while maintaining accurate physical interactions, facilitating benchmarking and dataset generation for vision-based robotics applications.
-The system supports complex environmental effects, such as smoke, fire, and water dynamics, which are critical for evaluating robotic performance under adverse conditions.
+The system supports complex environmental effects, such as smoke, fire, and water dynamics, which are critical to evaluating robotic performance under adverse conditions.
 We benchmark visual navigation and SLAM methods within our framework, demonstrating its utility for testing real-world robustness in controlled yet diverse scenarios.
 By bridging the gap between physics accuracy and photorealistic rendering, our framework provides a powerful tool for advancing robotics research and sim-to-real transfer.
+Our open-source framework is available at https://unrealroboticslab.github.io/.
 I INTRODUCTION
@@
 High-fidelity simulation environments provide realistic perception and physics interactions, bridging the gap between simulation and reality.
-However, achieving both photorealism and accurate physics simultaneously remains a challenge due to computational constraints and the complexity of real-world dynamics [1].
-This paper presents a novel simulation pipeline that combines the rendering capabilities of Unreal Engine [2] with the advanced physics simulation of MuJoCo [3] to create photorealistic environments for robotics applications.
+However, achieving both photorealism and accurate physics simultaneously remains a challenge due to computational constraints and the complexity of real-world dynamics [szot2021habitat].
+This paper presents a novel simulation pipeline that combines the rendering capabilities of Unreal Engine [unrealengine] with the advanced physics simulation of MuJoCo [todorov2012mujoco] to create photorealistic environments for robotics applications.
 By integrating these two frameworks, our approach delivers high-quality simulation results suitable for tasks such as perception training, dataset generation, and benchmarking.
-A key advantage of this pipeline is its ability to generate complex and dynamic environmental effects—such as smoke, fire, and water waves—which are crucial for training robots to handle diverse and unpredictable conditions.
-These effects enable the simulation of rare events that are difficult to capture in real-world datasets or reproduce in existing simulators.
+A key advantage of this pipeline is its ability to generate complex and dynamic environmental effects — such as smoke, fire, and water waves — which are crucial to training robots to handle diverse and unpredictable conditions.
+These effects enable the simulation of rare events that are difficult to capture in real-world datasets or reproduce in existing general-purpose robotics simulators.
 As a result, they can be used to benchmark vision-based policies under out-of-distribution conditions and generate training data for visual-based navigation.
-We present a novel, open-source simulation framework - the Unreal Robotics Lab (URL) that integrates Unreal Engine for photorealistic rendering with MuJoCo for accurate physics simulation.
-Unlike existing simulators focused on specific domains (e.g., autonomous driving or aerial robotics) or embodied-ai for limited robots, our system is designed for general-purpose robotic applications, supporting quadrupeds, manipulators, UAVs, and more.
-Additionally, we incorporate CoACD-based convex decomposition [4] for improved non-convex object representation.
+We present a novel, open-source simulation framework — the Unreal Robotics Lab (URL) that integrates Unreal Engine for photorealistic rendering with MuJoCo for accurate physics simulation.
+Unlike existing simulators focused on specific domains (e.g.
+autonomous driving or aerial robotics) or embodied AI for limited robots, our system is designed for general-purpose robotic applications, supporting quadrupeds, manipulators, UAVs, and more.
+While powerful general-purpose simulators such as Isaac Sim [NVIDIA_Isaac_Sim] offer broad capabilities, our framework prioritizes physical fidelity by utilizing the MuJoCo physics engine, which is highly validated for its accuracy and stability in contact-rich control tasks.
+Additionally, we incorporate CoACD-based convex decomposition [wei2022coacd] for improved non-convex object representation.
 Our key contributions are:
@@
   • Generalized Benchmarking for Diverse Conditions: A structured benchmarking system evaluates robotic perception and control under rare conditions (e.g., smoke, fog, dynamic lighting, dynamic movement).
-  • Scene Generation: A GPT-based module automates diverse environment creation, reducing manual setup and enabling adaptive scene variations.
-  • Comprehensive Sensor and Control Framework: Built-in support for cameras, IMUs, force sensors, and controllers for various robot types, with full Robot Operating System (ROS) compatibility.
+  • Comprehensive Sensor and Control Framework: Built-in support for cameras, IMUs, force sensors, and controllers for various robot types, with Robot Operating System (ROS) compatibility.
   • Open-Source and Extensible: The system will be released as open-source, supporting modular extensions for new controllers, planners, and robotic platforms.
-The subsequent sections are structured as follows.
-Sec.
-II provides an overview of the relevant literature, discussing previous works in the field, while Sec.
-III outlines the proposed method in detail.
-Sec.
-IV presents the experimental setup, including both simulation and real robot experiments.
-Finally, in Sec.
-V, we present the conclusions and discuss future
 II RELATED WORK
@@
 This section reviews related work in three main areas: robotics-focused simulation platforms, comparisons of physics engines, and benchmarking SLAM and navigation under adverse visual conditions.
-Several other game engines and simulation platforms offer photorealistic rendering [5, 6, 7], but Unreal Engine provides distinct advantages through Lumen global illumination and Nanite virtualized geometry.
-Unlike other simulators that rely on hardware-intensive ray tracing, Lumen [8] enables real-time global illumination and reflections without requiring precomputed lighting, making it well-suited for dynamic environments where lighting conditions change.
-Furthermore, Nanite’s automatic level of detail (LOD) [9] system allows highly detailed assets with millions of triangles to be rendered efficiently, without affecting performance.
+Several other game engines and simulation platforms offer photorealistic rendering [mittal2023orbit, puig2023habitat, dosovitskiy2017carla], but Unreal Engine provides distinct advantages through Lumen global illumination and Nanite virtualized geometry.
+Unlike other simulators that rely on hardware-intensive ray tracing, Lumen [unrealengine] enables real-time global illumination and reflections without requiring precomputed lighting, making it well-suited for dynamic environments where lighting conditions change.
+Furthermore, Nanite’s automatic level of detail (LOD) [unrealengine] system allows highly detailed assets with millions of triangles to be rendered efficiently, without affecting performance.
 This is particularly beneficial for large-scale or high-fidelity indoor environments where maintaining real-time performance is critical.
@@
 Several simulation platforms have been developed for robotic perception, control, and reinforcement learning, each with different trade-offs in realism and efficiency.
-Gazebo [10] is a widely-used open-source simulator that supports multiple physics engines; ODE [10], Simbody [11], (Py)Bullet [12], and DART [13], although its rendering capabilities lag behind modern game-engine-based simulators.
-MuJoCo [3] excels in efficient and accurate physics modeling for robotic control but lacks built-in photorealistic rendering.
-GPU-accelerated platforms such as Isaac Orbit [5] and Isaac Gym [14] enable large-scale reinforcement learning, prioritizing speed over advanced rendering.
-Intel SPEAR [15] provides photorealistic indoor environments suitable for embodied AI research.
-CARLA [7] and AirSim [16] leverage Unreal Engine for photorealistic urban and aerial simulations, with a primary focus on autonomous driving and drone applications.
+Gazebo [koenig2004design] is a widely-used open-source simulator that supports multiple physics engines (ODE [koenig2004design], Simbody [sherman2011simbody], (Py)Bullet [coumans2016pybullet], DART [lee2018dart]), although its rendering capabilities lag behind modern game-engine-based simulators.
+MuJoCo [todorov2012mujoco] excels in efficient and accurate physics modeling for robotic control but lacks built-in photorealistic rendering.
+GPU-accelerated platforms such as Isaac Orbit [mittal2023orbit] and Isaac Gym [makoviychuk2021isaac] enable large-scale reinforcement learning, prioritizing speed over advanced rendering.
+Isaac Sim [NVIDIA_Isaac_Sim], built on NVIDIA Omniverse, provides extensible robotics simulation with high-fidelity physics via PhysX [nvidia_physx]—which excels in angular stability and scalability but may lag in accuracy for contact dynamics and linear stability compared to MuJoCo.
+Intel SPEAR [simulator_intelspear] provides photorealistic indoor environments suitable for embodied AI research.
+CARLA [dosovitskiy2017carla] and AirSim [shah2018airsim] leverage Unreal Engine for photorealistic urban and aerial simulations, with a focus on autonomous driving and drone applications.
 CARLA, in particular, emphasizes high-level decision-making, such as traffic light interactions and multi-agent traffic flow, rather than fine-grained vehicle dynamics like tire skidding.
-However, both simulators remain largely domain-specific, limiting their utility for broader robotics applications beyond driving and aerial scenarios.
-RoboSuite [17] and RoboCasa [18] emphasize robotic manipulation using MuJoCo for precise physics but offer limited environmental complexity.
-MetaUrban [19] simulates realistic urban environments with dynamic traffic, beneficial for multi-agent interactions.
+However, both simulators are largely domain-specific, limiting their utility for broader robotics applications beyond driving and aerial scenarios.
+RoboSuite [robosuite2020] and RoboCasa [robocasa2024] emphasize robotic manipulation using MuJoCo for precise physics but offer limited environmental complexity.
 Although most simulators prioritize either physics accuracy or visual realism, few achieve both.
-Our approach bridges this gap by integrating MuJoCo for precise physics with Unreal Engine for state-of-the-art rendering, creating a high-fidelity simulation pipeline for diverse robotic applications.
+Our approach bridges this gap by integrating MuJoCo (precise physics) with Unreal Engine (state-of-the-art rendering), creating a high-fidelity simulation pipeline for diverse robotic applications.
 II-B Physics Engines for Robotics Simulation
 Physics engines are essential for realistic robotic simulation, particularly for modeling contact, soft-body dynamics, and fluid interactions.
-Various engines have been evaluated in terms of accuracy, efficiency, and scalability, with notable differences in their suitability for robotics [20].
-MuJoCo [3] excels in soft-contact modeling, accuracy, and stability, making it ideal for complex robotic tasks such as grasping and locomotion.
+Various engines have been evaluated in terms of accuracy, efficiency, and scalability, with notable differences in their suitability for robotics [erez2015simulation].
+MuJoCo [todorov2012mujoco] excels in soft-contact modeling, accuracy, and stability, making it ideal for complex robotic tasks such as grasping and locomotion.
 However, it lacks built-in high-fidelity rendering.
-ODE (Open Dynamics Engine) [10], commonly used in Gazebo, faces stability issues in high-DOF tasks and requires extensive parameter tuning.
-Bullet [12] is popular for reinforcement learning and manipulation, but its precision lags behind MuJoCo.
-Unreal Engine’s Chaos and Havok [21] are optimized for game physics and real-time effects, excelling in destruction, fluid, and particle-based simulations.
+ODE (Open Dynamics Engine) [koenig2004design], commonly used in Gazebo, faces stability issues in high-DOF tasks and requires extensive parameter tuning.
+Bullet [coumans2016pybullet] is popular for reinforcement learning and manipulation, but its precision lags behind MuJoCo.
+Unreal Engine’s Chaos and Havok physics solvers [havok_physics] are optimized for game physics and real-time effects, excelling in destruction, fluid, and particle-based simulations.
 However, they are not designed for robotic applications that need soft-contact modeling or high-precision joint dynamics.
-NVIDIA PhysX [22] offers GPU-accelerated simulation suitable for fast reinforcement learning, but its omission of Coriolis forces makes it unsuitable for high-accuracy control.
-Our approach combines MuJoCo’s precise physics with Unreal Engine’s advanced rendering, achieving both photorealism and accurate simulation for robotic applications.
+NVIDIA PhysX [nvidia_physx] offers GPU-accelerated simulation suitable for fast reinforcement learning, but its omission of Coriolis forces makes it unsuitable for high-accuracy control.
+Our approach combines MuJoCo’s precise physics, which are validated for their accuracy and stability [erez2015simulation], with Unreal Engine’s advanced rendering.
+This integration achieves both photorealism and accurate simulation for robotic applications.
+It also provides a broad feature set by inheriting the built-in capabilities of both Unreal Engine and MuJoCo.
 II-C Benchmarking SLAM and Navigation Under Adverse Visual Conditions
 Evaluating SLAM and navigation algorithms in diverse environments is challenging, as many existing datasets capture structured scenarios under ideal conditions, making it difficult to test performance under adverse effects such as smoke, fog or dynamic lighting.
-Envodat [23] provides a dataset tailored for challenging conditions, but real-world data collection remains limited in scale and diversity.
+Envodat [nwankwo2024envodat] provides a dataset tailored for challenging conditions, but real-world data collection remains limited in scale and diversity.
 Our simulation pipeline fills this gap by synthetically generating rare-event scenarios, enabling benchmarking and training of vision-based navigation policies in controlled, repeatable settings.
 Current simulation platforms often emphasize either high-precision physics (e.g., MuJoCo, Isaac Gym) or photorealistic rendering (e.g., Unreal-based simulators such as Carla and AirSim).
-A recent simulator, SPEAR [15], integrates MuJoCo with Unreal Engine but primarily focuses on providing a Python-based OpenAI Gym environment, pre-built scenes, limited robot support, and task-specific benchmarks.
-In contrast, our approach integrates MuJoCo’s precise physics with Unreal Engine’s advanced rendering in a flexible and generalized simulation environment.
-Our framework allows users to freely configure robots and tasks, and utilizes ROS to treat simulated interactions as closely as possible to real-world scenarios.
+While our work shares a similar goal with recent simulators such as SPEAR [simulator_intelspear] simulator in combining Unreal Engine with MuJoCo, our frameworks are built on fundamentally different design philosophies.
+SPEAR is primarily structured as a Python-based OpenAI Gym environment, making it well-suited for task-specific, algorithmic research in embodied AI.
+In contrast, our framework is designed for systems-level validation through its native ROS integration, allowing researchers to test a complete robotics software stack as it would be deployed in the real world.
+This focus on flexible, real-world system replication is further reflected in our integrated workflow.
+Our framework allows users to build and simulate scenes directly within the Unreal Editor, automatically handling all asset processing.
+This streamlined approach eliminates the need for separate scripts and manual data handling for importing and exporting assets, which can be required in other simulators.
+Ultimately, this design makes URL a generalized platform where users can freely import any MuJoCo-compatible robot and construct custom tasks, rather than focusing on a curated set of pre-built scenes and benchmarks.
 Furthermore, by leveraging Unreal’s Niagara Particle and Fluid effects, we enable simulation of adverse environmental conditions such as smoke, fire, and water waves, extending traditional benchmarking capabilities and providing robust tools for evaluating robotic perception and navigation under challenging scenarios.
+Beyond rendering, UE’s mature ecosystem accelerates robotics research.
+We utilize its built-in NavMesh and Behavior Tree tools to script complex, dynamic scenarios with moving NPCs and vehicles, creating environments that are not just visually realistic but also behaviorally rich.
+This is critical for testing planner robustness in non-static worlds, a feature often lacking in physics-focused simulators.
 III Simulator setup
 III-A Overview
-Our simulator (URL) integrates MuJoCo and Unreal Engine to provide a comprehensive simulation environment that closely mimics real-world interactions.
-Robots can be effortlessly integrated by adding corresponding MuJoCo XML descriptions and associated Unreal Engine assets.
-The system supports standard robot configurations and sensors, including cameras, IMUs, and force sensors, ensuring wide compatibility.
-The overall system can be seen in Figure 2.
-The system is comprised of both External components and the internal simulator.
-The system has a SimManager that is modular and manages submodules comprised of a Benchmark, Replay, and SceneGen modules.
-It manages all references to the UE data and is developed as a plugin for UE that can be added to any project.
-For communication, our simulator leverages ROS to enable seamless interaction with existing robotics frameworks.
-Furthermore, faster custom communication methods are being explored through ZeroMQ [24] for improved real-time performance.
-We treat the simulation environment as a surrogate real-world, facilitating the testing and validation of robotic behaviors and perception systems in realistic scenarios.
-III-B Physics Manager
-Our system integrates MuJoCo and CoACD [4] within the Unreal Engine (UE) for accurate physics simulation and mesh decomposition.
-These systems are currently compiled into the UE project, and extend a custom interface.
-To represent objects in the scene, we introduce an intermediate representation, termed PhysicsObject, which encapsulates both geometric and state information.
-This allows us to efficiently manage and transform scene elements between UE and MuJoCo while ensuring that transformations and dynamic interactions are correctly synchronized.
-Objects within the scene are categorized into predefined types.
-Specifically, assets are tagged under one of the following categories: EP_STATIC, EP_STATIC_COMPLEX, EP_DYNAMIC, EP_DYNAMIC_COMPLEX, and EP_PRIMITIVE_(TYPE).
-Static objects remain unchanged throughout the simulation, whereas dynamic objects are actively tracked, with their transformations updated asynchronously within UE.
-The distinction between simple and complex dynamic objects determines whether the object undergoes convex decomposition by CoACD.
-CoACD decomposes meshes into a series of convex hulls, whereas the regular mujoco compile will generate only one convex hull per object.
-This allows CoACD to represent non-convex geometry and is recommended by MuJoCo [20].
-This classification is recorded within the corresponding PhysicsObject to guide further processing.
+Our simulator is implemented as a UE plugin centered around a custom SimManager Actor, with the overall architecture shown in Fig.
+2.
+This Actor manages several core sub-systems: a SceneManager for processing assets and constructing the MuJoCo scene, a PhysicsManager to synchronize MuJoCo physics with UE rendering, and Benchmark and Replay systems for performance evaluation and data logging.
+The pipeline begins with the SceneManager scanning for actors with a UCustomPhysicsComponent, processing their properties to construct a MuJoCo Specification (MjSpec), and compiling the final physics model.
+The system supports standard robots defined via MuJoCo XML with corresponding UE assets, including sensors such as cameras, IMUs, and force sensors.
+Communication leverages a modified ROS plugin [mania19scenarios] for compatibility with existing robotics frameworks, allowing the simulator to act as a high-fidelity, real-world surrogate for testing robotics behaviors and perception.
+To complement the physics and data pipeline, our framework uses the larger set of features of Unreal Engine to build rich and dynamic worlds.
+This includes high-fidelity rendering for perception research (Lumen, Nanite), rapid world-building with automated NavMesh generation for pathfinding, and complex scenario creation using Behavior Trees for intelligent agents and the Niagara system for adverse weather and effects.
+III-B Scene Building and Asset Processing
+UE actors are configured for MuJoCo using a UCustomPhysicsComponent, which contains an intermediate PhysicsObject to synchronize geometry and state between the two engines.
+The settings of this component dictate how the object is processed for simulation.
+Its geometric complexity setting (Complex or Simple) determines how meshes are converted into convex hulls for MuJoCo’s collision system, using either CoACD [wei2022coacd] for decomposition or direct compiler approximation.
+The component also controls the object’s mobility (static or dynamic) via an override or by defaulting to the actor’s native UE setting.
 Figure 1: Unreal Engine Landscape conversion to MuJoCo HeightField.
-Once asset classification is complete, mesh-based assets are exported as OBJ files, ensuring compatibility with MuJoCo.
-Additionally, if landscapes are present in the scene, they are scanned and converted into height-fields to be used within the simulation.
-This process uses a custom collision channel on the landscape, then, utilizing an autogenerated bounding box, we perform multithreaded raycasts to determine the heights to use.
-This bounding-box limits the raycasts to within this area only at a predefined resolution.
-This prevents the whole landscape from being scanned.
-Figure 1 demonstrates the output of this process.
-This method was chosen because landscapes can be very large, and the user may not be interested in converting the entire landscape for physics.
-It also has the added benefit of being able to model small clutter and decorations in the heightfield itself.
-This process enables the seamless reconstruction of large-scale terrains within MuJoCo, preserving critical environmental features necessary for accurate physics interactions.
+The SceneManger then iterates over all actors with the UCustomPhysicsComponent and processes the assets with their containing parameters.
+Meshes that are processed with CoACD are saved in the project’s SavedDir for caching.
+UE landscapes are converted to MuJoCo height fields using multithreaded raycasts within a user-defined bounding box at a set resolution, as shown in Fig.
+1.
+This targeted approach avoids processing entire large-scale terrains while still capturing fine details such as small clutter, preserving key environmental features for the physics simulation.
 Figure 2: Unreal - MuJoCo Simulator system diagram
-Following asset processing, an XML scene description is automatically generated to initialize MuJoCo.
-This XML encapsulates all relevant object data, including mesh file references, physics properties, and hierarchical relationships.
-Special attention is paid to robotic assets that are detected within the MuJoCo environment and mapped back to their corresponding UE entities.
-During this process, cameras and sensors specified in the XML are instantiated within UE to ensure consistency across both environments.
-Sensor elements are represented as sites within MuJoCo, allowing for precise placement and interaction within the physics simulation.
+MjSpec is then used to generate the final MuJoCo Model and XML file.
+Robotic assets, packaged as UE pawns with their own XML, are automatically detected and added into the MjSpec.
+Sensors are defined as sites in the robot’s XML, and we employ a hybrid sensing approach: non-visual data (IMUs, force sensors) is sourced directly from MuJoCo, while visual data (cameras, LiDAR) is rendered in UE for photorealism.
+The exported XML and assets can also be used to run the scene as a standalone MuJoCo simulation, independent of UE.
+III-C Physics Manager
 The physics simulation operates asynchronously, running within a separate execution thread to maximize performance.
-Transformation updates are also managed in a dedicated asynchronous thread within UE to maintain real-time synchronization with MuJoCo.
+Specifically, the MuJoCo physics state is advanced by calling mj_step within a dedicated thread that maintains a configurable physics timestep (e.g., 1000 Hz) by yielding control between steps.
+Meanwhile, the Unreal Engine rendering thread operates independently at its own rate, typically 60 FPS or higher, depending on scene complexity.
+Transformation updates are managed in a dedicated asynchronous thread within UE to maintain real-time synchronization with MuJoCo.
+To ensure data integrity, locking mechanisms are used when the rendering thread accesses the latest physics state, a design that mimics the behavior of MuJoCo’s native simulate GUI and decouples rendering performance from physics fidelity.
 External forces applied within MuJoCo can be propagated to UE entities, allowing for bidirectional interaction between the simulation and the rendered environment.
 UE physics can still be used for objects not managed by MuJoCo.
-To improve computational efficiency, assets that remain unchanged between simulation runs are cached, preventing unnecessary recomputation and redundant file exports.
-This caching mechanism ensures that static elements do not incur additional processing overhead, allowing the system to focus on dynamic entities that require continuous state updates.
-The final XML scene description, along with the automatically generated mesh assets, enables direct execution within MuJoCo as a standalone simulation.
-A supporting script is provided to facilitate this process, allowing the same environment to be used independently of UE when required.
-III-C Benchmark System
-The benchmarking system automates the recording and management of evaluation metrics throughout the simulation.
-At its core, it consists of a MetricManager and a BaseMetric class, providing a structured and extensible framework to track performance indicators.
-The BaseMetric class defines a fundamental interface for all metrics, encapsulating a two-dimensional float array that serves as a storage mechanism for time-series data.
-This class provides essential methods, including initialization, resetting, and exporting data to CSV format.
-These methods ensure that each metric implementation adheres to a consistent lifecycle, simplifying the process of integrating new metrics into the system.
-Derived metrics can override these base methods to implement domain-specific logic while still benefiting from the structured data storage and export mechanisms.
-The MetricManager is responsible for the automatic discovery and registration of all metric implementations.
-It dynamically populates a list of available metrics by scanning for corresponding class definitions, allowing modular and scalable metric tracking.
-Users can enable or disable specific metrics through a dedicated user interface, ensuring flexibility in selecting relevant performance indicators for a given experiment.
-By default, the system tracks core metrics, including: DistanceToGoal, which tracks the remaining distance from an entity to a predefined target location; Collisions, which records contact events with obstacles or other entities; TimeToGoal, which measures the time required to reach the destination; and GlobalPose, which logs the absolute position and orientation of an entity over time.
-All metrics are recorded at discrete intervals and automatically saved as CSV files within designated experiment folders, ensuring organized and accessible results.
-The system also features grace periods (delaying data recording for $n$ seconds) and timeouts (terminating recordings after $n$ seconds), improving efficiency and reproducibility across simulation runs.
-III-D Scene Generation
-Figure 3: Example of room generated using the experimental Scene Generation system leveraging GPT API.
-Our simulator features an experimental scene generation module designed primarily for indoor environments, leveraging OpenAI’s ChatGPT [25] API.
-When an instance of the ASceneGenerator Actor is present and assigned a RoomType, the system automatically generates object placements based on the specified scene type (Fig.
-3).
-The module queries ChatGPT with a structured prompt containing a list of available assets and receives a JSON-formatted response specifying object positions.
-These placements are then parsed, and the corresponding assets are instantiated in the environment.
-To ensure spatial coherence, the system detects overlapping objects using Unreal Engine’s physics system.
-If overlaps occur, an iterative refinement loop provides feedback to ChatGPT for updated placement recommendations, continuing until all collisions are resolved or a predefined iteration limit is reached.
-This method enables efficient, collision-free scene generation, producing diverse and structured environments suitable for robotic experimentation.
-III-E Replay
-Our system features an automated recording and replay mechanism for sensor and body pose data, enabling efficient data collection and experimentation.
-It uses custom structs to track only essential information –pose data, sensor readings, and timestamps– storing themin a queue for minimal overhead and precise sequence reproduction.
-Data streams are managed through separate JSON files, all of which can be dynamically saved and loaded at runtime.
+III-D Communication
+Our framework uses a modified ROS plugin [mania19scenarios] that automates the communication layer.
+This plugin automatically publishes all sensor data from both MuJoCo and UE to their corresponding ROS topics, the names of which are derived directly from the sensor definitions in the robot’s XML file.
+Similarly, for robot control, the plugin dynamically generates subscribers for any defined actuators or joints.
+This allows standard robotics controllers to send actuator or joint commands directly to MuJoCo.
+III-E Benchmark System
+The benchmarking system automates metric recording and management through a structured, extensible framework built upon a MetricManager and a BaseMetric class.
+The BaseMetric class provides a fundamental interface for all metrics, handling time-series data storage and defining a consistent lifecycle with methods for initialization, reset, and CSV export.
+New metrics can be created by deriving from this class to implement domain-specific logic while inheriting the core data-handling functionality.
+The MetricManager automatically discovers and registers all metric classes by scanning project definitions, allowing users to select and configure them for each experiment via a UI.
+By default, the system tracks core metrics such as DistanceToGoal (distance to target), Collisions (contact events), TimeToGoal (task completion time), and GlobalPose (absolute pose).
+All metrics are recorded at discrete intervals and saved as CSV files in designated experiment folders, with support for configurable grace periods and timeouts to ensure reproducible results.
+III-F Replay System
+Our system has an automated recording and replay mechanism for sensor and body pose data, allowing efficient data collection and experimentation.
+It uses custom structs to track only essential information, pose data, sensor readings, and timestamps, storing them in a queue for minimal overhead and precise sequence reproduction.
+Data streams are managed through separate JSON files that can be dynamically saved and loaded at runtime.
 By publishing replayed data on the same ROS topics as live data, the system integrates seamlessly with existing pipelines.
-This allows researchers to collect data once and replay it in visually modified environments, supporting domain adaptation, robustness testing, and dataset augmentation for machine learning and benchmarking.
-III-F Extended Features
-Our system leverages Unreal Engine’s advanced rendering and simulation capabilities, surpassing traditional robotics simulators like MuJoCo [3], Gazebo [10], and Isaac Sim [5].
-Key features include real-time global illumination (Lumen), high-detail mesh rendering (Nanite), and physically based reflections, which enhance perception-based robotics research.
-Procedural animations, NavMesh navigation, and built-in environmental tools enable efficient creation of dynamic simulations, while the Niagara particle system facilitates testing under adverse visual conditions (e.g., smoke, fog).
-Additionally, packaged standalone deployments improve reproducibility and scalability, addressing limitations of conventional simulators.
+This allows researchers to collect data once and replay them in visually modified environments, supporting domain adaptation, robustness testing, and dataset enhancement for machine learning and benchmarking.
 III-G Example Deployment of Robots, Controllers, and Visual Planners
-The system provides a unified framework where robots, controllers, and planners operate as in real-world deployments.
+The system provides a unified framework in which robots, controllers, and planners operate as in real-world deployments.
 All components communicate through ROS, ensuring seamless integration with external systems.
@@
 III-G2 Controllers
-The system includes joint controllers based on position and torque, as well as more advanced controllers such as Walk-These-Ways [26] and VBC [27].
-UAVs use a PID-based controller, whereas manipulators rely on inverse kinematics (IK) methods such as Mink [28].
+The system includes joint controllers based on position and torque, as well as more advanced controllers such as Walk-These-Ways [margolis2023walk] and VBC [liu2024visual].
+UAVs use a PID-based controller, whereas manipulators rely on inverse kinematics (IK) methods such as Mink [Zakka_Mink_Python_inverse_2024].
 Since controllers operate within the ROS ecosystem, new control strategies can be easily integrated.
 III-G3 Planners
-Motion planning is handled using learning-based and optimization-based planners such as ViNT [29], GNM [30], and NoMAD [31].
+Motion planning is handled using learning-based and optimization-based planners such as ViNT [shah2023vint], GNM [shah2023gnm], and NoMAD [sridhar2024nomad].
 The ROS-based architecture of the framework ensures compatibility with additional external planners.
 III-G4 Visual SLAM
-The system includes visual SLAM techniques such as OrbSLAM2 [32], OrbSLAM3 [33], and MASt3R-SLAM [34], enabling real-time localization and mapping.
+The system includes visual SLAM techniques such as OrbSLAM2 [murORB2], OrbSLAM3 [campos2021orb], and MASt3R-SLAM [murai2024mast3r], enabling real-time localization and mapping.
 These methods integrate directly into the framework, enabling realistic perception and mapping in simulated environments.
@@
 Lower values indicate closer alignment.
-To compare models trained on real-world data, we use CLIP [35] and the EfficientNet-B0 [36] visual encoder used in ViNT [29].
+To compare models trained on real-world data, we use CLIP [radford2021learning] and the EfficientNet-B0 [tan2019efficientnet] visual encoder used in ViNT [shah2023vint].
 We first capture an image from the real world and recreate it within our simulator.
@@
 The similarity in attention patterns also suggests that key objects remain distinguishable across domains, which is critical for transfer learning and sim-to-real applications.
-To further quantify similarity, we compute KL divergence (0.4679) and cosine similarity (0.7157) between real and simulated feature distributions.
-Higher cosine similarity and lower KL divergence suggest closer alignment.
-While we did not aim to replicate real-world scenes with identical assets, opting instead for a simulation that remains visually representative of real conditions.
-Although the assets in our simulation differ from those in the real environment, our findings, together with GradCAM and EigenCAM comparisons, indicate that the simulation is visually representative enough for our evaluation purposes.
+We further quantified the visual similarity between our hand-crafted simulations and real-world scenes by computing the KL divergence and cosine similarity for three scenes.
+These simulations were not built as photo-realistic replicas but as a proof of concept that our system can generate scenes visually representative of real conditions.
+As detailed in Table I, the Hallway scene, for example, yielded a KL divergence of 0.4679 and a cosine similarity of 0.7157, and an average of 0.4739 and 0.7094 in all scenes.
+These findings, along with our GradCAM and EigenCAM comparisons, confirm that our simulation is visually realistic enough for our evaluation purposes.
+Metric Office Hallway Flat Hallway Bedroom Avg
+KL divergence 0.4679 0.4435 0.5103 0.4739
+Cosine similarity 0.7157 0.7231 0.6895 0.7094
+TABLE I: Image comparison metrics between real-world scenes and simulated recreations.
 IV-B Visual Navigation and SLAM Benchmarking
@@
 We benchmark standard methods to demonstrate simulator realism.
-We evaluated visual navigation (Experiment 1) with GNM [30] and ViNT [29], and visual SLAM (Experiment 2) with OrbSLAM2 [32], OrbSLAM3 [33], and MASt3R-SLAM [34].
+We evaluated visual navigation (Experiment 1) with GNM [shah2023gnm] and ViNT [shah2023vint], and visual SLAM (Experiment 2) with ORB-SLAM2 [murORB2], ORB-SLAM3 [campos2021orb], and MASt3R-SLAM [murai2024mast3r].
 Each method was run 5 times per condition, averaging results across trials.
@@
 The warehouse and residential house were used for this experiment.
-Methods were not fine-tuned, relying on a simplified topological map approach from the authors’ original repository [31], which may limit performance [37].
-Success Weighted by Collision (SC). Following [38], we adopt the SC metric:
+Methods were not fine-tuned, relying on a simplified topological map approach from the authors’ original repository [sridhar2024nomad], which may limit performance [visualnav-transformer-issue5].
+Success Weighted by Collision (SC). Following [eftekhar2024one], we adopt the SC metric:
 $SC=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{1}{1+c_{i}},$ (5)
@@
 Success
-$\uparrow$
+[…]
 […]
@@
 ViNT 0.8 4.4 90.14 2.59 0.3667
-TABLE I: Performance of different Visual navigation methods under various environments and varying adverse effects.
-Table I presents the results, including Weighted SC to factor in collision penalties.
+TABLE II: Performance of different Visual navigation methods under various environments and varying adverse effects.
+Table II presents the results, including Weighted SC to factor in collision penalties.
 State-of-the-art methods perform well under normal conditions, with ViNT generally outperforming GNM.
@@
 In severe adversity, neither method succeeds in the Warehouse, leading to zero SC.
+These failures are likely attributable to the dense, volumetric smoke simulated by the Niagara system, which caused persistent occlusion and aliasing of key visual features, preventing the vision-based planners from identifying a valid path.
 This suggests that these models struggle in extreme conditions due to limited training data, highlighting the need for training in diverse high-adversity conditions, which can be generated more effectively in simulation.
@@
 $\text{Coverage}=\frac{\sum_{i=1}^{n-1}\left\|\mathbf{p}_{i+1}-\mathbf{p}_{i}\right\|}{\sum_{i=1}^{n-1}\left\|\hat{\mathbf{p}}_{i+1}-\hat{\mathbf{p}}_{i}\right\|}$ (7)
-note that since monocular SLAM estimates trajectories with arbitrary scale, Umeyama’s method [39] is applied for scale and pose correction before computing coverage.
+Note that since monocular SLAM estimates trajectories with arbitrary scale, Umeyama’s method [umeyama1991least] is applied for scale and pose correction before computing coverage.
 Poor tracking or alignment can result in coverage exceeding 100%.
 Scaled ATE adjusts ATE for localization failures:
-Scaled $\text{Scaled ATE}=\frac{\text{ATE}}{\text{Coverage}}$ (8)
+$\text{Scaled ATE}=\frac{\text{ATE}}{\text{Coverage}}$ (8)
 Figure 5: comparision of different Visual SLAM algorithms under different visual effects
@@
 Cover-
-$\uparrow$
+[…]
 […]
@@
 Cover-
-$\uparrow$
+[…]
 […]
@@
 Cover-
-$\uparrow$
+[…]
 Ware- house None 4.23 6.61 0.639 5.29 7.72 0.685 0.324 0.322 1.01
@@
 Major 0.475 2.53 0.188 2.70 11.6 0.233 0.275 0.284 0.966
-TABLE II: Performance of Different VSLAM Methods under Varying Visual Effects.
-From Table II and Fig.
-6, Visual SLAM performance degrades with increasing visual adversity, highlighting sensitivity to challenging conditions.
-OrbSLAM2 struggles the most, with sharp declines in coverage and increased ATE, particularly in the warehouse environment.
-OrbSLAM3 performs well in clean conditions but deteriorates significantly under adversity, suggesting reduced robustness to occlusions, lighting changes, and dynamic elements.
+TABLE III: Performance of Different VSLAM Methods under Varying Visual Effects.
+From Table III and Fig.
+5, Visual SLAM performance degrades with increasing visual adversity, highlighting sensitivity to challenging conditions.
+ORB-SLAM2 struggles the most, with sharp declines in coverage and increased ATE, particularly in the warehouse environment.
+ORB-SLAM3 performs well in clean conditions but deteriorates significantly under adversity, suggesting reduced robustness to occlusions, lighting changes, and dynamic elements.
 MASt3R-SLAM maintains the lowest ATE and highest coverage across conditions, demonstrating superior resilience to visual disturbances.
@@
 These replay-based evaluations provide a controlled, repeatable framework for assessing SLAM robustness across different levels of adversity, reinforcing the need for training and evaluation in diverse, high-adversity conditions to enhance real-world deployment reliability.
+IV-B3 Experiment 3: Real-world
+Visual SLAM algorithms were evaluated on realworld tunnel dataset [kubelka2024we] under two conditions: clear and smoky, where a section of the tunnel was obscured by smoke generated from a smoke machine.
+Examples of these conditions are shown in Fig.
+6.
+The percentage of tracked frames under these conditions was measured for various visual SLAM algorithms, results can be seen in Tab.
+IV, where it can be seen that the tracking performance of all SLAM algorithm degrades significantly under smoky conditions, consistent with the findings from our simulated experiments.
+Env Adver. Frames tracked $\%$$\uparrow$
+OrbSLAM2
+Mono
+OrbSLAM3
+Mono
+Mast3R
+SLAM
+Mine Clean 7.83 59.6 100
+Smoky 0.00 38.5 43.6
+TABLE IV: Performance of Different VSLAM Methods under Varying Visual Effects in the Real-World.
+Figure 6: Examples of Realworld Tunnel data [kubelka2024we] used for evaluating VSLAM algorithm under adverse visual conditions.
+IV-B4 Discussion
+Our experiments collectively demonstrate that current state-of-the-art visual navigation and SLAM algorithms, while proficient in ideal conditions, remain brittle to the types of realistic, adverse visual phenomena that are produced by our framework and that are encountered in the realworld.
+The observed performance degradation highlights a critical gap in current training and testing paradigms.
+Futhermore, because adverse real-world datasets are scarce, there is a reinforced need for high-fidelity simulators such as URL, which can generate controlled, repeatable, and diverse high-adversity scenarios to advance the development of more robust robotic perception systems.
 V CONCLUSION
 This paper presents a high-fidelity robotics simulation framework that integrates advanced photorealistic rendering with precise physics modeling to support vision-based robotics research and benchmarking.
-By leveraging Unreal Engine’s rendering capabilities and MuJoCo’s accurate physics, our system enables evaluation of robotic navigation and SLAM methods under diverse and adverse environmental conditions.
+Using Unreal Engine’s rendering capabilities and MuJoCo’s accurate physics, our system enables evaluation of robotic navigation and SLAM methods under diverse and adverse environmental conditions.
 Experimental results confirm that our simulator provides a robust testbed for evaluating real-world robustness in controlled, repeatable scenarios.
 Despite its strengths, the system has certain limitations, including the absence of direct reinforcement learning (RL) support, lack of deformable terrain simulation, limited scene generation capabilities, and a focus on general robotics rather than autonomous vehicles or high-precision drone simulations.
-Future Work.
-To address these limitations and further enhance our simulator, we plan to further enhance the MuJoCo integration, allowing the simulator to communicate with a standalone MuJoCo process and run RL environments (MuJoCo Playground [40]), generate synthetic datasets to enhance vision-based model training, expand automated scene generation, and enable imitation learning via VR-based teleoperation.
+To address these limitations and further enhance our simulator, we plan to further enhance the MuJoCo integration, allowing the simulator to communicate with a standalone MuJoCo process and run RL environments (MuJoCo Playground [zakka2025mujocoplayground]), generate synthetic datasets to enhance vision-based model training, expand automated scene generation, and enable imitation learning via VR-based teleoperation.
 Overall, our platform serves as a robust foundation for benchmarking and data generation, bridging the gap between simulation and real-world robotics deployment.

### v3

@@
 I INTRODUCTION
-Simulation plays a crucial role in robotics research, enabling safe, efficient, and cost-effective development, testing, and validation of robotic systems before real-world deployment.
-High-fidelity simulation environments provide realistic perception and physics interactions, bridging the gap between simulation and reality.
-However, achieving both photorealism and accurate physics simultaneously remains a challenge due to computational constraints and the complexity of real-world dynamics [szot2021habitat].
-This paper presents a novel simulation pipeline that combines the rendering capabilities of Unreal Engine [unrealengine] with the advanced physics simulation of MuJoCo [todorov2012mujoco] to create photorealistic environments for robotics applications.
-By integrating these two frameworks, our approach delivers high-quality simulation results suitable for tasks such as perception training, dataset generation, and benchmarking.
-A key advantage of this pipeline is its ability to generate complex and dynamic environmental effects — such as smoke, fire, and water waves — which are crucial to training robots to handle diverse and unpredictable conditions.
-These effects enable the simulation of rare events that are difficult to capture in real-world datasets or reproduce in existing general-purpose robotics simulators.
-As a result, they can be used to benchmark vision-based policies under out-of-distribution conditions and generate training data for visual-based navigation.
-We present a novel, open-source simulation framework — the Unreal Robotics Lab (URL) that integrates Unreal Engine for photorealistic rendering with MuJoCo for accurate physics simulation.
-Unlike existing simulators focused on specific domains (e.g.
-autonomous driving or aerial robotics) or embodied AI for limited robots, our system is designed for general-purpose robotic applications, supporting quadrupeds, manipulators, UAVs, and more.
-While powerful general-purpose simulators such as Isaac Sim [NVIDIA_Isaac_Sim] offer broad capabilities, our framework prioritizes physical fidelity by utilizing the MuJoCo physics engine, which is highly validated for its accuracy and stability in contact-rich control tasks.
-Additionally, we incorporate CoACD-based convex decomposition [wei2022coacd] for improved non-convex object representation.
+Simulation is essential for robotics research, enabling safe, efficient development, testing, and validation of robotic systems before real-world deployment.
+High-fidelity environments must simultaneously deliver realistic perception and accurate physics interactions—an open challenge due to competing computational demands and the complexity of real-world dynamics [szot2021habitat].
+We introduce the Unreal Robotics Lab (URL), a novel open-source framework that tightly integrates Unreal Engine’s state-of-the-art rendering pipeline (Lumen global illumination, Nanite virtualized geometry, Niagara particles) [unrealengine] with MuJoCo’s validated high-precision physics engine [todorov2012mujoco].
+By embedding MuJoCo natively inside the Unreal ecosystem via shared-memory execution, URL achieves deterministic physics while exposing Unreal’s full authoring, environmental-effect, and tools directly to roboticists.
+A key capability is the effortless generation of complex dynamic phenomena — smoke, fire, water waves, dynamic lighting, and moving agents—that are rare or hazardous to capture in real datasets.
+These effects enable controlled benchmarking of vision-based policies under out-of-distribution conditions and the creation of large-scale training data for perception and navigation.
+Unlike domain-specific simulators (e.g., autonomous driving or aerial robotics) or platforms focused on limited embodied-AI scenarios, URL is designed for general-purpose robotic applications, supporting quadrupeds, mobile manipulators, UAVs, and arms.
+While simulators such as Isaac Sim [NVIDIA_Isaac_Sim] offer broad capabilities, URL prioritises physical fidelity by retaining MuJoCo’s superior accuracy and stability in contact-rich tasks and incorporates CoACD-based convex decomposition [wei2022coacd] for improved non-convex object representation.
 Our key contributions are:
-  • High-Fidelity Physics with Photorealistic Rendering: Synchronization of MuJoCo’s physics with Unreal Engine’s advanced rendering (Lumen, Nanite, Niagara) enables realistic perception-based training and benchmarking.
-  • Generalized Benchmarking for Diverse Conditions: A structured benchmarking system evaluates robotic perception and control under rare conditions (e.g., smoke, fog, dynamic lighting, dynamic movement).
-  • Comprehensive Sensor and Control Framework: Built-in support for cameras, IMUs, force sensors, and controllers for various robot types, with Robot Operating System (ROS) compatibility.
-  • Open-Source and Extensible: The system will be released as open-source, supporting modular extensions for new controllers, planners, and robotic platforms.
+  • Native “Unreal-First” Architecture: We introduce a shared-memory execution model that embeds the MuJoCo engine directly within the Unreal Engine environment.
+This eliminates the latency and synchronization overhead of traditional client-server models.
+  • High-Fidelity Physics-to-Visual Synchronization: By mirroring the MuJoCo schema within the UE Actor-Component hierarchy, we enable seamless synchronization between MuJoCo’s stable physics and UE’s photorealistic rendering pipeline (Lumen, Nanite, Niagara).
+  • Automated Environmental Processing and Geometry Conversion: The framework provides a robust pipeline for bridging complex UE environmental assets to physics-ready formats, utilizing multi-threaded terrain sampling and automated CoACD-based conversion for arbitrary static meshes.
+  • Comprehensive ROS-Compatible Middleware: A unified interface for sensors and actuators that supports diverse robotic platforms with native support for ROS and ZeroMQ communication.
+This enables the “plug-and-play” deployment of modern planners (e.g., ViNT, NoMAD) and SLAM frameworks (e.g., ORB-SLAM3) directly within the simulation.
+  • Generalized Benchmarking for Diverse Conditions: We provide a structured framework for evaluating robotic systems under rare or extreme environmental conditions, complemented by a deterministic replay system that allows for robust domain randomization testing across varied visual environments.
 II RELATED WORK
-Robotic simulation has advanced significantly in recent years, with various frameworks addressing key challenges such as physics accuracy, photorealistic rendering, and real-time performance.
-This section reviews related work in three main areas: robotics-focused simulation platforms, comparisons of physics engines, and benchmarking SLAM and navigation under adverse visual conditions.
-Several other game engines and simulation platforms offer photorealistic rendering [mittal2023orbit, puig2023habitat, dosovitskiy2017carla], but Unreal Engine provides distinct advantages through Lumen global illumination and Nanite virtualized geometry.
-Unlike other simulators that rely on hardware-intensive ray tracing, Lumen [unrealengine] enables real-time global illumination and reflections without requiring precomputed lighting, making it well-suited for dynamic environments where lighting conditions change.
-Furthermore, Nanite’s automatic level of detail (LOD) [unrealengine] system allows highly detailed assets with millions of triangles to be rendered efficiently, without affecting performance.
-This is particularly beneficial for large-scale or high-fidelity indoor environments where maintaining real-time performance is critical.
-Beyond rendering, Unreal Engine offers a flexible and extensible ecosystem, supporting third-party plugins, visual scripting, AI-driven NPCs, and cross-platform deployment, making it highly adaptable for a variety of robotic simulation tasks.
-These features ensure that simulations can achieve both high physical accuracy and photorealistic perception, making them an ideal choice for vision-based robotics research.
+Robotic simulation has advanced significantly in recent years, addressing key challenges in physics accuracy, photorealistic rendering, and real-time performance.
+This section reviews related work in three main areas: robotics-focused simulation platforms, physics engine comparisons, and benchmarking of SLAM and navigation under adverse visual conditions.
+Game engines and simulation platforms offer photorealistic rendering [mittal2023orbit, puig2023habitat, dosovitskiy2017carla], yet Unreal Engine provides distinct advantages through Lumen global illumination and Nanite virtualized geometry [unrealengine].
+Lumen enables real-time global illumination and reflections without precomputed lighting—ideal for dynamic environments—while Nanite efficiently renders highly detailed assets with millions of triangles.
+Combined with its extensible ecosystem (third-party plugins, visual scripting, AI-driven NPCs, and cross-platform deployment), Unreal Engine is particularly well-suited for vision-based robotics research.
 II-A Robotics-Focused Simulation Platforms
-Robotic simulation has made considerable progress, with numerous frameworks addressing challenges in physics accuracy, rendering quality, and real-time performance.
-This section reviews related work in three primary areas: robotics simulation platforms, comparisons of physics engines, and benchmarking of SLAM and navigation under adverse visual conditions.
-Several simulation platforms have been developed for robotic perception, control, and reinforcement learning, each with different trade-offs in realism and efficiency.
-Gazebo [koenig2004design] is a widely-used open-source simulator that supports multiple physics engines (ODE [koenig2004design], Simbody [sherman2011simbody], (Py)Bullet [coumans2016pybullet], DART [lee2018dart]), although its rendering capabilities lag behind modern game-engine-based simulators.
-MuJoCo [todorov2012mujoco] excels in efficient and accurate physics modeling for robotic control but lacks built-in photorealistic rendering.
-GPU-accelerated platforms such as Isaac Orbit [mittal2023orbit] and Isaac Gym [makoviychuk2021isaac] enable large-scale reinforcement learning, prioritizing speed over advanced rendering.
-Isaac Sim [NVIDIA_Isaac_Sim], built on NVIDIA Omniverse, provides extensible robotics simulation with high-fidelity physics via PhysX [nvidia_physx]—which excels in angular stability and scalability but may lag in accuracy for contact dynamics and linear stability compared to MuJoCo.
-Intel SPEAR [simulator_intelspear] provides photorealistic indoor environments suitable for embodied AI research.
-CARLA [dosovitskiy2017carla] and AirSim [shah2018airsim] leverage Unreal Engine for photorealistic urban and aerial simulations, with a focus on autonomous driving and drone applications.
-CARLA, in particular, emphasizes high-level decision-making, such as traffic light interactions and multi-agent traffic flow, rather than fine-grained vehicle dynamics like tire skidding.
-However, both simulators are largely domain-specific, limiting their utility for broader robotics applications beyond driving and aerial scenarios.
-RoboSuite [robosuite2020] and RoboCasa [robocasa2024] emphasize robotic manipulation using MuJoCo for precise physics but offer limited environmental complexity.
-Although most simulators prioritize either physics accuracy or visual realism, few achieve both.
-Our approach bridges this gap by integrating MuJoCo (precise physics) with Unreal Engine (state-of-the-art rendering), creating a high-fidelity simulation pipeline for diverse robotic applications.
+Several simulators target robotic perception, control, and reinforcement learning [ellis2022navigation, shanks2025dreamernav], each with different trade-offs in realism and efficiency.
+Gazebo [koenig2004design] is a widely-used open-source platform that supports multiple physics engines (ODE [koenig2004design], Simbody [sherman2011simbody], Bullet [coumans2016pybullet], DART [lee2018dart]), although its rendering lags behind modern game-engine solutions.
+MuJoCo [todorov2012mujoco] excels in accurate, efficient physics for control tasks but lacks built-in photorealistic rendering.
+GPU-accelerated frameworks such as Isaac Orbit [mittal2023orbit] and Isaac Gym [makoviychuk2021isaac] prioritise speed for large-scale reinforcement learning, while Isaac Sim [NVIDIA_Isaac_Sim] (built on NVIDIA Omniverse) uses PhysX [nvidia_physx] for scalability at the expense of contact-dynamics accuracy compared with MuJoCo.
+Intel SPEAR [simulator_intelspear] provides photorealistic indoor environments for embodied AI, and domain-specific simulators such as CARLA [dosovitskiy2017carla] and AirSim [shah2018airsim] leverage Unreal Engine for urban/aerial scenarios.
+Manipulation-focused suites RoboSuite [robosuite2020] and RoboCasa [robocasa2024] rely on MuJoCo physics but offer limited environmental complexity.
+Although most platforms emphasise either physics fidelity or visual realism, few achieve both.
+URL bridges this gap by natively integrating MuJoCo’s precise physics with Unreal Engine’s state-of-the-art rendering in an editor-first, ROS-native framework designed for general-purpose robotics (quadrupeds, mobile manipulators, UAVs, and arms).
 II-B Physics Engines for Robotics Simulation
-Physics engines are essential for realistic robotic simulation, particularly for modeling contact, soft-body dynamics, and fluid interactions.
-Various engines have been evaluated in terms of accuracy, efficiency, and scalability, with notable differences in their suitability for robotics [erez2015simulation].
-MuJoCo [todorov2012mujoco] excels in soft-contact modeling, accuracy, and stability, making it ideal for complex robotic tasks such as grasping and locomotion.
-However, it lacks built-in high-fidelity rendering.
-ODE (Open Dynamics Engine) [koenig2004design], commonly used in Gazebo, faces stability issues in high-DOF tasks and requires extensive parameter tuning.
-Bullet [coumans2016pybullet] is popular for reinforcement learning and manipulation, but its precision lags behind MuJoCo.
-Unreal Engine’s Chaos and Havok physics solvers [havok_physics] are optimized for game physics and real-time effects, excelling in destruction, fluid, and particle-based simulations.
-However, they are not designed for robotic applications that need soft-contact modeling or high-precision joint dynamics.
-NVIDIA PhysX [nvidia_physx] offers GPU-accelerated simulation suitable for fast reinforcement learning, but its omission of Coriolis forces makes it unsuitable for high-accuracy control.
-Our approach combines MuJoCo’s precise physics, which are validated for their accuracy and stability [erez2015simulation], with Unreal Engine’s advanced rendering.
-This integration achieves both photorealism and accurate simulation for robotic applications.
-It also provides a broad feature set by inheriting the built-in capabilities of both Unreal Engine and MuJoCo.
+Physics engines are critical for realistic contact, soft-body, and fluid modelling.
+MuJoCo [todorov2012mujoco] is widely favored for its for soft-contact accuracy and stability in complex robotic tasks [erez2015simulation].
+ODE (used in Gazebo) suffers from stability issues in high-DOF scenarios and requires extensive tuning.
+Bullet [coumans2016pybullet] is popular for reinforcement learning and manipulation but lags MuJoCo in precision.
+NVIDIA PhysX [nvidia_physx] offers GPU acceleration suitable for fast RL yet omits Coriolis forces and exhibits lower linear-stability accuracy.
+Unreal Engine’s Chaos and Havok solvers [havok_physics] excel at real-time destruction and particle effects but are not optimised for high-precision joint dynamics or soft-contact robotics.
+The continued adoption of MuJoCo’s physics is further evidenced by the fact that next-generation GPU-accelerated simulators such as Genesis [Genesis] and Newton [newton_physics] have adopted MJWarp [mujoco_warp] (a high-performance GPU backend directly derived from MuJoCo) as their core physics engine.
+This widespread use in the latest state-of-the-art platforms confirms that MuJoCo’s contact modelling, accuracy, and stability remain unmatched—even as the field shifts toward massive parallel GPU simulation.
+URL therefore retains native MuJoCo as the sole physics backend while inheriting Unreal Engine’s advanced rendering, particle systems, and tools—delivering both photorealism and validated physical fidelity without compromise.
 II-C Benchmarking SLAM and Navigation Under Adverse Visual Conditions
-Evaluating SLAM and navigation algorithms in diverse environments is challenging, as many existing datasets capture structured scenarios under ideal conditions, making it difficult to test performance under adverse effects such as smoke, fog or dynamic lighting.
-Envodat [nwankwo2024envodat] provides a dataset tailored for challenging conditions, but real-world data collection remains limited in scale and diversity.
-Our simulation pipeline fills this gap by synthetically generating rare-event scenarios, enabling benchmarking and training of vision-based navigation policies in controlled, repeatable settings.
-Current simulation platforms often emphasize either high-precision physics (e.g., MuJoCo, Isaac Gym) or photorealistic rendering (e.g., Unreal-based simulators such as Carla and AirSim).
-While our work shares a similar goal with recent simulators such as SPEAR [simulator_intelspear] simulator in combining Unreal Engine with MuJoCo, our frameworks are built on fundamentally different design philosophies.
-SPEAR is primarily structured as a Python-based OpenAI Gym environment, making it well-suited for task-specific, algorithmic research in embodied AI.
-In contrast, our framework is designed for systems-level validation through its native ROS integration, allowing researchers to test a complete robotics software stack as it would be deployed in the real world.
-This focus on flexible, real-world system replication is further reflected in our integrated workflow.
-Our framework allows users to build and simulate scenes directly within the Unreal Editor, automatically handling all asset processing.
-This streamlined approach eliminates the need for separate scripts and manual data handling for importing and exporting assets, which can be required in other simulators.
-Ultimately, this design makes URL a generalized platform where users can freely import any MuJoCo-compatible robot and construct custom tasks, rather than focusing on a curated set of pre-built scenes and benchmarks.
-Furthermore, by leveraging Unreal’s Niagara Particle and Fluid effects, we enable simulation of adverse environmental conditions such as smoke, fire, and water waves, extending traditional benchmarking capabilities and providing robust tools for evaluating robotic perception and navigation under challenging scenarios.
-Beyond rendering, UE’s mature ecosystem accelerates robotics research.
-We utilize its built-in NavMesh and Behavior Tree tools to script complex, dynamic scenarios with moving NPCs and vehicles, creating environments that are not just visually realistic but also behaviorally rich.
-This is critical for testing planner robustness in non-static worlds, a feature often lacking in physics-focused simulators.
-III Simulator setup
-III-A Overview
-Our simulator is implemented as a UE plugin centered around a custom SimManager Actor, with the overall architecture shown in Fig.
-2.
-This Actor manages several core sub-systems: a SceneManager for processing assets and constructing the MuJoCo scene, a PhysicsManager to synchronize MuJoCo physics with UE rendering, and Benchmark and Replay systems for performance evaluation and data logging.
-The pipeline begins with the SceneManager scanning for actors with a UCustomPhysicsComponent, processing their properties to construct a MuJoCo Specification (MjSpec), and compiling the final physics model.
-The system supports standard robots defined via MuJoCo XML with corresponding UE assets, including sensors such as cameras, IMUs, and force sensors.
-Communication leverages a modified ROS plugin [mania19scenarios] for compatibility with existing robotics frameworks, allowing the simulator to act as a high-fidelity, real-world surrogate for testing robotics behaviors and perception.
-To complement the physics and data pipeline, our framework uses the larger set of features of Unreal Engine to build rich and dynamic worlds.
-This includes high-fidelity rendering for perception research (Lumen, Nanite), rapid world-building with automated NavMesh generation for pathfinding, and complex scenario creation using Behavior Trees for intelligent agents and the Niagara system for adverse weather and effects.
-III-B Scene Building and Asset Processing
-UE actors are configured for MuJoCo using a UCustomPhysicsComponent, which contains an intermediate PhysicsObject to synchronize geometry and state between the two engines.
-The settings of this component dictate how the object is processed for simulation.
-Its geometric complexity setting (Complex or Simple) determines how meshes are converted into convex hulls for MuJoCo’s collision system, using either CoACD [wei2022coacd] for decomposition or direct compiler approximation.
-The component also controls the object’s mobility (static or dynamic) via an override or by defaulting to the actor’s native UE setting.
+Evaluating SLAM and navigation algorithms under adverse conditions (smoke, fog, dynamic lighting, or moving agents) remains difficult, as most public datasets and simulators capture only structured, well-lit scenes [EmbleyRiches2025SAGESA].
+Envodat [nwankwo2024envodat] provides one challenging real-world dataset, yet scale and diversity are still limited.
+URL addresses this limitation by synthetically generating controllable rare-event scenarios via Niagara particles and fluids, together with dynamic NPCs and vehicles scripted through Unreal’s NavMesh and Behavior Trees.
+Current simulators typically prioritise either high-precision physics (e.g., MuJoCo-based) or photorealistic rendering (e.g., CARLA, AirSim).
+While our work shares the high-level goal of SPEAR [simulator_intelspear] in combining Unreal rendering with MuJoCo physics, the design philosophies differ fundamentally.
+SPEAR is structured as a Python-based OpenAI Gym environment for task-specific embodied-AI research.
+In contrast, URL is built for systems-level validation: it offers native ROS integration, direct scene authoring inside the Unreal Editor (with automatic asset processing and CoACD convex decomposition), and seamless import of any MuJoCo-compatible robot.
+This enables full-stack testing of complete robotic software pipelines exactly as they would be deployed in the real world, together with repeatable benchmarking of perception and planning robustness in behaviorally rich, non-static environments.
+III Simulator Architecture
+III-A System Overview
+We propose an “Unreal Engine-first” simulation framework (Fig.
+1) that embeds the MuJoCo physics engine entirely within the Unreal Engine (UE) ecosystem.
+By discarding traditional client-server communication models in favor of in-process, shared-memory execution, the architecture ensures absolute physical determinism.
+This design enables the physics simulation to natively leverage UE’s photorealistic rendering pipeline (Lumen, Nanite), environmental effects (Niagara), and AI navigation (NavMesh) without the latency or synchronization overhead inherent in distributed architectures.
+Figure 1: Unreal - MuJoCo Simulator system diagram.
+III-B World Authoring and Asset Instantiation
+A core principle of this framework is the comprehensive mirroring of the MuJoCo schema within UE’s Actor-Component hierarchy.
+Rather than providing an abstracted wrapper, the system exposes nearly the entire MuJoCo specification—including global defaults, solver options, and runtime configurations—as native, strongly-typed properties directly within the UE interface.
+The framework caters to both legacy MuJoCo users and native Unreal Engine development through a tiered ingestion pipeline.
+Users can integrate legacy models by dragging standard MuJoCo XML files directly into the UE Content Browser.
+This triggers a dedicated factory that parses the MJCF specification to automatically generate a structurally equivalent UE Blueprint.
+Articulated robots and environmental objects are instantiated as AMjArticulation Actors, which manage a recursive tree of localized components such as UMjBody, UMjGeom, UMjJoint, and UMjTendon.
+This process automatically imports and converts all referenced meshes and primitives into native UE assets, creating a persistent representation that benefits from UE’s advanced material and rendering systems.
+For native authoring, users can instantiate an empty AMjArticulation and procedurally build the hierarchy through the standard UE “Add” interface.
+The architecture abstracts away from ID-based indexing and manual pointer arithmetic; users can access live physics data directly through the articulation’s component references without managing raw mjModel or mjData addresses.
+Scene construction is finalized by placing these Blueprints within a level containing an AMjManager Actor.
+Upon initialization, the manager performs a scene-wide scan to aggregate all active Actors into a contiguous specification.
+To optimize performance, the framework utilizes asynchronous asset serialization and caching for mesh generation, ensuring complex geometries are processed once.
+The system maintains bidirectional compatibility, allowing users to export UE-authored levels back to standardized MuJoCo formats.
+III-C Environmental Sampling and Geometry Processing
+To bridge the gap between UE’s diverse environmental assets and MuJoCo’s collision requirements, the framework provides automated conversion utilities.
+The AMjHeightfieldActor facilitates terrain interaction by executing multi-threaded, bounded raycasts to sample the environment.
+This actor is not restricted to UE Landscapes; it can sample any geometry not already managed by the MuJoCo solver, specifically any Actor lacking an AMjArticulation or a UMjQuickConvertComponent.
+This allows for the generation of optimized MuJoCo heightfields from complex, non-articulated environmental geometry while minimizing memory overhead (Fig.
+2).
+For discrete objects, the UMjQuickConvertComponent enables the instant conversion of arbitrary UE static meshes into MuJoCo-compatible collision bodies using CoACD [wei2022coacd] for automatic convex hull decomposition.
 Figure 2: Unreal Engine Landscape conversion to MuJoCo HeightField.
-The SceneManger then iterates over all actors with the UCustomPhysicsComponent and processes the assets with their containing parameters.
-Meshes that are processed with CoACD are saved in the project’s SavedDir for caching.
-UE landscapes are converted to MuJoCo height fields using multithreaded raycasts within a user-defined bounding box at a set resolution, as shown in Fig.
-1.
-This targeted approach avoids processing entire large-scale terrains while still capturing fine details such as small clutter, preserving key environmental features for the physics simulation.
-Figure 2: Unreal - MuJoCo Simulator system diagram
-MjSpec is then used to generate the final MuJoCo Model and XML file.
-Robotic assets, packaged as UE pawns with their own XML, are automatically detected and added into the MjSpec.
-Sensors are defined as sites in the robot’s XML, and we employ a hybrid sensing approach: non-visual data (IMUs, force sensors) is sourced directly from MuJoCo, while visual data (cameras, LiDAR) is rendered in UE for photorealism.
-The exported XML and assets can also be used to run the scene as a standalone MuJoCo simulation, independent of UE.
-III-C Physics Manager
-The physics simulation operates asynchronously, running within a separate execution thread to maximize performance.
-Specifically, the MuJoCo physics state is advanced by calling mj_step within a dedicated thread that maintains a configurable physics timestep (e.g., 1000 Hz) by yielding control between steps.
-Meanwhile, the Unreal Engine rendering thread operates independently at its own rate, typically 60 FPS or higher, depending on scene complexity.
-Transformation updates are managed in a dedicated asynchronous thread within UE to maintain real-time synchronization with MuJoCo.
-To ensure data integrity, locking mechanisms are used when the rendering thread accesses the latest physics state, a design that mimics the behavior of MuJoCo’s native simulate GUI and decouples rendering performance from physics fidelity.
-External forces applied within MuJoCo can be propagated to UE entities, allowing for bidirectional interaction between the simulation and the rendered environment.
-UE physics can still be used for objects not managed by MuJoCo.
-III-D Communication
-Our framework uses a modified ROS plugin [mania19scenarios] that automates the communication layer.
-This plugin automatically publishes all sensor data from both MuJoCo and UE to their corresponding ROS topics, the names of which are derived directly from the sensor definitions in the robot’s XML file.
-Similarly, for robot control, the plugin dynamically generates subscribers for any defined actuators or joints.
-This allows standard robotics controllers to send actuator or joint commands directly to MuJoCo.
-III-E Benchmark System
-The benchmarking system automates metric recording and management through a structured, extensible framework built upon a MetricManager and a BaseMetric class.
-The BaseMetric class provides a fundamental interface for all metrics, handling time-series data storage and defining a consistent lifecycle with methods for initialization, reset, and CSV export.
-New metrics can be created by deriving from this class to implement domain-specific logic while inheriting the core data-handling functionality.
-The MetricManager automatically discovers and registers all metric classes by scanning project definitions, allowing users to select and configure them for each experiment via a UI.
-By default, the system tracks core metrics such as DistanceToGoal (distance to target), Collisions (contact events), TimeToGoal (task completion time), and GlobalPose (absolute pose).
-All metrics are recorded at discrete intervals and saved as CSV files in designated experiment folders, with support for configurable grace periods and timeouts to ensure reproducible results.
-III-F Replay System
-Our system has an automated recording and replay mechanism for sensor and body pose data, allowing efficient data collection and experimentation.
-It uses custom structs to track only essential information, pose data, sensor readings, and timestamps, storing them in a queue for minimal overhead and precise sequence reproduction.
-Data streams are managed through separate JSON files that can be dynamically saved and loaded at runtime.
-By publishing replayed data on the same ROS topics as live data, the system integrates seamlessly with existing pipelines.
-This allows researchers to collect data once and replay them in visually modified environments, supporting domain adaptation, robustness testing, and dataset enhancement for machine learning and benchmarking.
-III-G Example Deployment of Robots, Controllers, and Visual Planners
+III-D Unified Interaction and Bi-Directional Physics
+To maintain strict determinism, the MuJoCo simulation executes on a dedicated asynchronous thread.
+The solver steps at the specific timestep defined within the MuJoCo model, which is exposed and configurable via the AMjManager.
+Interaction is unified across the framework’s API; every MuJoCo property, function, and sensor is accessible via both C++ and UE Blueprints.
+This includes a comprehensive sensor interface where all MuJoCo-native sensors are exposed as UE components, providing researchers with Blueprint nodes to access real-time data streams.
+Furthermore, the framework includes a native graphical UI (MjSimulateWidget) that replicates the functionality of MuJoCo’s standalone Simulate, allowing for real-time visualization, state monitoring, and interactive control within the UE viewport.
+While the framework handles pointer management, users retain the ability to interact with the raw MuJoCo structs if required for custom solver logic.
+This flexibility is complemented by a robust one-way kinematic coupling (mocap) which allows UE’s native Chaos physics and AI systems to influence the MuJoCo solver in real-time.
+For example, an NPC governed by the UE NavMesh can be assigned a UMjQuickConvertComponent with its “Unreal Driven” property enabled, transforming it into a mocap body that moves in synchronization with UE’s physics.
+III-E Communication, Metrics, and Replay
+The framework supports multiple communication modalities, allowing users to choose between a modified ROS plugin [mania19scenarios] or a high-performance ZeroMQ (ZMQ) bridge.
+These plugins automate the mapping of MuJoCo and UE sensors to their respective topics or sockets based on the robot’s configuration.
+For empirical evaluation, a MetricManager provides a structured framework for recording time-series data directly to CSV.
+Furthermore, a high-fidelity Replay System caches joint kinematics and sensor streams, enabling researchers to record physical trajectories and deterministically replay them across iteratively augmented visual environments for robust domain randomization testing.
+III-F Example Deployment of Robots, Controllers, and Visual Planners
 The system provides a unified framework in which robots, controllers, and planners operate as in real-world deployments.
 All components communicate through ROS, ensuring seamless integration with external systems.
-III-G1 Example Robot Integrations
+III-F1 Example Robot Integrations
 The framework supports quadrupeds, quadruped mobile manipulators, UAVs, and robotic manipulators, enabling diverse experimental setups for locomotion, aerial navigation, and manipulation.
-III-G2 Controllers
+III-F2 Controllers
 The system includes joint controllers based on position and torque, as well as more advanced controllers such as Walk-These-Ways [margolis2023walk] and VBC [liu2024visual].
@@
 Since controllers operate within the ROS ecosystem, new control strategies can be easily integrated.
-III-G3 Planners
+III-F3 Planners
 Motion planning is handled using learning-based and optimization-based planners such as ViNT [shah2023vint], GNM [shah2023gnm], and NoMAD [sridhar2024nomad].
 The ROS-based architecture of the framework ensures compatibility with additional external planners.
-III-G4 Visual SLAM
+III-F4 Visual SLAM
 The system includes visual SLAM techniques such as OrbSLAM2 [murORB2], OrbSLAM3 [campos2021orb], and MASt3R-SLAM [murai2024mast3r], enabling real-time localization and mapping.
@@
 IV-A Simulation and Real-World Comparision
-To evaluate the alignment between simulated and real-world visual inputs for image encoders, we employ Grad-CAM, EigenCAM, cosine similarity and KL Divergence.
+To evaluate the alignment between simulated and real-world visual inputs for image encoders, we employ Grad-CAM [selvaraju2017grad], EigenCAM [muhammad2020eigen], cosine similarity and KL Divergence [kullback1951information].
 These methods assess both spatial attention and feature distribution similarity.
-Grad-CAM computes the gradient of the class score $y^{c}$ with respect to feature map activations:
+Specifically, spatial attention is computed using Grad-CAM and EigenCAM in equations 1 and 2 respectively to assess feature distribution similarity.
+Grad-CAM (Gradient-weighted Class Activation Mapping) computes gradients of class score $y^{c}$ (pre-softmax logit for class $c$) w.r.t.
+feature maps $A^{k}$ (spatial activations from channel $k$ in the final conv layer):
 $\alpha_{k}^{c}=\frac{1}{Z}\sum_{i}\sum_{j}\frac{\partial y^{c}}{\partial A^{k}_{ij}},\quad L^{c}=\text{ReLU}\left(\sum_{k}\alpha_{k}^{c}A^{k}\right)$ (1)
+where $\alpha_{k}^{c}$ is the Global Average Pooling (GAP)-weighted importance of channel $k$ for class $c$ ($i,j$: spatial indices; $Z$: width $\times$ height).
+Heatmap $L^{c}$ is the weighted sum of maps with ReLU, yielding a saliency map of regions influencing the model’s class $c$ prediction.
 EigenCAM generates class-agnostic saliency maps via principal component analysis (PCA):
 $C=\frac{1}{Z}\sum_{i,j}(A^{k}_{ij}-\mu)(A^{k}_{ij}-\mu)^{T},\quad L_{\text{EigenCAM}}=v_{1}A^{k}$ (2)
-where $\mu$ is the mean activation and $v_{1}$ is the first principal component.
+where $A_{ij}$ is the activation vector across channels at position $i,j$; $\mu$ is its mean over spatial locations, and $v_{1}$ is the first principal component (eigenvector of $C$).
+The map $L$ is the projection onto $v_{1}$.
 Figure 3: Comparison of real-world (left) and simulated (right) images, along with their corresponding Grad-CAM (CLIP) and EigenCAM (EfficientNet-B0) heatmaps. The highlighted class label is ”cone”.
@@
 Using Unreal Engine’s rendering capabilities and MuJoCo’s accurate physics, our system enables evaluation of robotic navigation and SLAM methods under diverse and adverse environmental conditions.
-Experimental results confirm that our simulator provides a robust testbed for evaluating real-world robustness in controlled, repeatable scenarios.
-Despite its strengths, the system has certain limitations, including the absence of direct reinforcement learning (RL) support, lack of deformable terrain simulation, limited scene generation capabilities, and a focus on general robotics rather than autonomous vehicles or high-precision drone simulations.
-To address these limitations and further enhance our simulator, we plan to further enhance the MuJoCo integration, allowing the simulator to communicate with a standalone MuJoCo process and run RL environments (MuJoCo Playground [zakka2025mujocoplayground]), generate synthetic datasets to enhance vision-based model training, expand automated scene generation, and enable imitation learning via VR-based teleoperation.
+Experimental results, such as utilizing the Replay System to test varied visual effects on identical trajectories, and Niagara-driven adversity benchmarks highlighting planner failures in volumetric smoke, confirm that our simulator provides a robust testbed for evaluating real-world robustness in controlled, repeatable scenarios.
+Despite its strengths, the system has certain limitations, including the lack of deformable terrain simulation, one-way kinematic coupling for UE driven Actors, and the absence of validated sensor noise models for depth and LiDAR data.
+To address these limitations and further improve our simulator, we plan to focus on narrowing the sim-to-real perceptual gap by implementing procedural material randomization, generate large-scale synthetic datasets to enhance vision-based model training, introduce VR teleop for data generation, expand automated scene generation, and enable bilateral physical interactions between MuJoCo and UE.
 Overall, our platform serves as a robust foundation for benchmarking and data generation, bridging the gap between simulation and real-world robotics deployment.